### v2

# Meta Co-Training: Two Views are Better than One

Abstract
In many practical computer vision scenarios unlabeled data is plentiful, but labels are scarce and difficult to obtain. As a result, semi-supervised learning which leverages unlabeled data to boost the performance of supervised classifiers have received significant attention in recent literature. One major class of semi-supervised algorithms is co-training. In co-training two different models leverage different independent and sufficient “views” of the data to jointly make better predictions. During co-training each model creates pseudo labels on unlabeled points which are used to improve the other model. We show that in the common case when independent views are not available we can construct such views inexpensively using pre-trained models. Co-training on the constructed views yields a performance improvement over any of the individual views we construct and performance comparable with recent approaches in semi-supervised learning, but has some undesirable properties. To alleviate the issues present with co-training we present Meta Co-Training which is an extension of the successful Meta Pseudo Labels approach to two views. Our method achieves new state-of-the-art performance on ImageNet-10% with very few training resources, as well as outperforming prior semi-supervised work on several other fine-grained image classification datasets.
1 Introduction
In many practical machine learning scenarios we have access to a large amount of unlabeled data, and relatively few labeled data points for the task on which we would like to train. For generic computer vision (CV) tasks there are large well-known and open source datasets with hundreds of millions or even billions of unlabeled images [62, 63]. In contrast, labeled data is usually scarce and expensive to obtain requiring many human-hours to generate. Hence, in this context, semi-supervised learning (SSL) methods are becoming increasingly popular as they rely on training useful learning models by using small amounts of labeled data and large amounts of unlabeled data.
Not unrelated to, and not to be confused with the idea of semi-supervised learning, is self-supervised learning. [^1] There is also an SSL technique called self-training. Recognizing the unfortunate similarity of these three terms, SSL will always mean semi-supervised learning in this text and we will not abbreviate the other terms. In self-supervised learning a model is trained with an objective that does not require a label that is not evident from the data itself. Self-supervised learning was popularized by the BERT [32] model for generating word embeddings for natural language processing (NLP) tasks. BERT generates its embeddings by solving the pretext task of masked word prediction. Pretext tasks present unsupervised objectives that are not directly related to any particular supervised objective we might desire to solve, but rather are solved in the hope of learning a suitable representation to more easily solve a downstream task. Due to the foundational nature of these models to the success of many models in NLP, or perhaps the fact that they serve as a foundation for other models to be built upon, these pretext learners are often referred to as foundation models. Inspired by foundation models from NLP, several pretext tasks have been proposed for CV with associated foundation models that have been trained to solve them [8, 38, 20, 42, 23, 49, 19, 58, 24, 43]. Unlike NLP, the learned representations for images are often much smaller than the images themselves which yields the additional benefit of reduced computational effort when using the learned representations.
Typically, SSL requires the learner to make predictions on a large amount of unlabeled data as well as the labeled data during training. Furthermore, SSL often requires that the learner be re-trained [12, 76, 50, 72, 48, 55, 39, 70, 74, 40, 17] partially, or from scratch. Given the computational benefits of learning from a frozen representation, combining SSL and self-supervised learning is a natural idea. Additionally, the representation learner allows us to compress our data to a smaller size. This compression is beneficial not only because of the reduction in operations required that is associated with a smaller model, but also because it allows us to occupy less space on disk, in RAM, and in high bandwidth memory on hardware accelerators.
These benefits are particularly pronounced for co-training [12] in which two different “views” of the data must be obtained or constructed and then two different classifiers must be trained and re-trained iteratively. Seldom does a standard problem in machine learning present two independent and sufficient views of the problem to be leveraged for co-training. As a result, if one hopes to apply co-training one usually needs to construct two views from the view that they have. To take advantage of the computational advantages of representation learning, and the large body of self-supervised literature, we propose using the embeddings produced by foundation models trained on different pretext tasks as the views for co-training. To our knowledge we are the first to propose using frozen learned feature representations as views for co-training classification, though we think the idea is very natural. We show that views constructed in this way appear to satisfy the necessary conditions for the effectiveness of co-training.
Unfortunately, while we can boost performance by training different models on two views, co-training performs sub-optimally during its pseudo-labeling step. To more effectively leverage two views we propose a novel method for co-training based on Meta Pseudo Labels [59] called Meta Co-Training in which each model is trained to provide better pseudo labels to the other given only its view of the labeled data and the performance of the other model on the labeled set. We show that our approach provides an improvement over both co-training and the current state-of-the-art (SotA) on the ImageNet-10% dataset, as well as it establishes new SotA few-shot performance on several other fine-grained image classification tasks.
Summary of Contributions.   Our contributions are therefore summarized as follows.
  1. 1.
    Use of embeddings obtained from different pre-trained models as different views for co-training.
  2. 2.
    We boost the performance of standard co-training by proposing a method that relies on Meta Pseudo Labels [59], called Meta Co-Training.
  3. 3.
    We establish new state-of-the-art top-1 accuracy on ImageNet-10% [31] and new state-of-the-art top-1 few-shot accuracy on several other CV classification tasks.
  4. 4.
    We make our implementation publicly available:
    https://github.com/JayRothenberger/Meta-Co-Training
Outline of the Paper.   In Section 2 we discuss related work. In Section 3 we establish notation and provide more information on co-training and meta pseudo labels. In Section 4 we describe our proposed methods of view construction and of meta co-training. In Section 5 we present and discuss our experimental findings on ImageNet [31], Flowers102 [57], Food101 [14], FGVCAircraft [54], iNaturalist [44], and iNaturalist 2021 [45]. In Section 6 we conclude with a summary and ideas for future work.
2 Related Work
While SSL [21, 78, 67, 73] has become very relevant in recent years with the influx of the data collected by various sensors, the paradigm and its usefulness have been studied at least since the 1960’s; e.g., [41, 29, 64, 36]. Among the main approaches for SSL that are not directly related to our work, we find techniques that try to identify the maximum likelihood parameters for mixture models [30, 7], constrained clustering [9, 28, 69, 68, 18, 3], graph-based methods [11, 10, 53, 6], as well as purely theoretical probably approximately correct (PAC)-style results [4, 37, 25]. Below we discuss SSL approaches closer to our work.
Self-Training and Student-Teacher Frameworks.   One of the earliest ideas for SSL is that of self-training [64, 36]. This idea is still very popular with different student-teacher frameworks [79, 76, 59, 50, 72, 66, 55, 39, 17]. This technique is also frequently referred to as pseudo-labeling [48] as the student tries to improve its performance by making predictions on the unlabeled data, then integrating the most confident predictions to the labeled set, and subsequently retraining using the broader labeled set. In addition, one can combine ideas; e.g., self-training with clustering [46], or create ensembles via self-training [47], potentially using graph-based information [53]. An issue with these approaches is that of confirmation bias [2], where the student cannot improve a lot due to incorrect labeling that occurs either because of self-training, or due to a fallible teacher.
Co-Training.   Another paradigmatic family of SSL algorithms are co-training [12] algorithms, where instances have two different views for the same underlying phenomenon. Co-training has been very successful with wide applicability [34] and can outperform learning algorithms that use only single-view data [56]. Furthermore, the generalization ability of co-training is well-motivated theoretically [12, 27, 5, 26]. There are several natural extensions of the base idea, such as methods for constructing different views [71, 15, 22, 65, 16], including connections to active learning [33], methods that are specific to deep learning [60], and other natural extensions that exploit more than two views on data; e.g., [77].
3 Background
We briefly describe two important components of semi-supervised learning that are used jointly in order to devise our proposed new technique. However, before we do so, we also define basic notions so that we can eliminate ambiguity.
3.1 Notation
Supervised learning is based on the idea of learning a function, called a model, $f$ belonging to some model space $\mathcal{F}$, using training examples $(x,y)$ that exemplify some underlying phenomenon. Training examples are labeled instances; i.e., $x\in\mathcal{X}$ is an instance drawn from some instance space $\mathcal{X}$ and $y$ is a label from some label set $\mathcal{Y}$. This relationship between $\mathcal{X}$ and $\mathcal{Y}$ can be modeled with a probability distribution $D$ that governs $\mathcal{X}\times\mathcal{Y}$. We write $(x,y)\sim D$ to indicate that a training example $(x,y)$ is drawn from $\mathcal{X}\times\mathcal{Y}$ according to $D$. We use $\mathbf{1}\left\{\mathcal{A}\right\}$ as an indicator function of the event $\mathcal{A}$; that is, $\mathbf{1}\left\{\mathcal{A}\right\}$ is equal to 1 when $\mathcal{A}$ holds, otherwise 0. For vectors $\alpha,\beta\in\mathbb{R}^{d}$ we indicate their Hadamard (element-wise) product using $\otimes$; that is, $\gamma=\alpha\otimes\beta$, where $\gamma_{j}=\alpha_{j}\cdot\beta_{j}$ for $j\in\{1,\ldots,d\}$.
As is standard in classification problems, we let model $f$ be a function of the form $f\colon\mathcal{X}\rightarrow\Delta^{|\mathcal{Y}|}$ where $\Delta^{|\mathcal{Y}|}$ is the $|\mathcal{Y}|$-dimensional unit simplex. Equivalently $f$ can be understood as a probability distribution $\mathbf{Pr}(y\ |\ x)$. We use $f(x)|_{j}$ to denote the $j$-th value of $f$ in the output in such a case. It is common for $\mathcal{F}$ to be a parameterized family of functions (e.g. artificial neural networks). We use $\theta$ as a vector that holds such parameters, and we denote the function parameterized by $\theta$ as $f_{\theta}$. An individual prediction $f_{\theta}(x)$ for an instance $x$ is evaluated via the use of a loss function $\ell$ that determines the quality of the prediction $f_{\theta}(x)$ with respect to the actual value $y$ that should have been predicted. For classification problems a natural way of defining loss functions is of the form $\ell\colon\mathcal{Y}\times\Delta^{|\mathcal{Y}|}\rightarrow\mathbb{R}_{+}$. For the entirety of this text $\ell$ will refer to the cross-entropy loss which is defined as $\ell\left(y,f_{\theta}(x)\right)=-\sum_{\xi\in\mathcal{Y}}\mathbf{1}\left\{\xi=y\right\}\log\left(f_{\theta}(x)|_{\xi}\right)$.
We measure the fitness of a model $f_{\theta}$ by its (true) risk $R_{D}\left(f_{\theta}\right)=\mathbf{E}_{(x,y)\sim D}\left[\ell\left(y,f_{\theta}(x)\right)\right]$. While we want to create a model $f_{\theta}$ that has low true risk, in principle we apply empirical risk minimization (ERM) which yields a model $f_{\theta}$ that minimizes the risk over a set of training examples. That is, we use empirical risk as a proxy for the true risk. Given a set $S=\left\{(x_{i},y_{i})\right\}_{i=1}^{\left\lvert S\right\rvert}$, we denote the empirical risk of $f_{\theta}$ on $S$ as $\widehat{R}_{S}\left(f_{\theta}\right)=\frac{1}{\left\lvert S\right\rvert}\sum_{i=1}^{\left\lvert S\right\rvert}\ell\left(y,f_{\theta}(x)\right)$. To ease notation, when the inputs to $\ell$ are a batch of examples $(X,Y)$, we denote $\widehat{R}_{(X,Y)}\left(f_{\theta}\right)$ as simply $\ell(Y,f_{\theta}(X))$. Similarly, when we write the arg max of a batch of vectors we refer to the batch of the arg max of the individual values. In general, ERM can be computationally hard (e.g., [13, 1]); hence, we often rely on gradient-based optimization techniques, such as stochastic gradient descent (SGD), that can take us to a local optimum efficiently. We use $\eta$ as the learning rate and $T$ as the maximum number of steps (or updates) in such a gradient-based optimization procedure.
Semi-Supervised Learning.   We are interested in situations where we have a large pool of unlabeled instances $U$ as well as a small portion $L$ of them that is labeled. In other words, we assume that the learning algorithm has access to a dataset $S=L\cup U$, such that $L=\left\{(x_{i},y_{i})\right\}_{i=1}^{m}$ and $U=\left\{x_{j}\right\}_{j=1}^{u}$. In this setting, semi-supervised (SSL) methods attempt to first learn an initial model $f_{init}$ using the labeled data $L$ via a supervised learning algorithm, and in sequence, an attempt is being made in order to harness the information that is hidden in the unlabeled set $U$, so that a better model $f$ can be learnt.
3.2 Meta Pseudo Labels
In order to discuss Meta Pseudo Labels, first we need to discuss Pseudo Labels. Pseudo Labels (PL) [48] is a broad paradigm for SSL, which involves a teacher and a student in the learning process. The teacher looking at the unlabeled set $U$, selects a few instances and assigns pseudo-labels to them. Subsequently these pseudo-examples are shown to the student as if they were legitimate examples so that the student can learn a better model. In this context, the teacher $f_{\theta_{T}}$ is fixed and the student is trying to learn a better model by aligning its predictions to pseudo-labeled batches $(X_{u},f_{\theta_{T}}(X_{u}))$. Thus, on a pseudo-labeled batch, the student optimizes its parameters using
$\theta_{S}^{\text{PL}}\in\operatorname*{arg\,min}_{\theta_{S}}\ell\left(\operatorname*{arg\,max}_{\xi}f_{\theta_{T}}(X_{u})|_{\xi},f_{\theta_{S}}(X_{u})\right)\,.$ (1)
The hope is that updated student parameters $\theta_{S}^{\text{PL}}$ will behave well on the labeled data $L=(X_{L},Y_{L})=\left\{(x_{i},y_{i})\right\}_{i=1}^{m}$; i.e., the loss $\mathcal{L}_{L}\left(\theta_{S}^{\text{PL}}\left(\theta_{T}\right)\right)=\ell\left(Y_{L},f_{\theta_{S}^{\text{PL}}}(X_{L})\right)$ is small, where we defined $\mathcal{L}\left(\theta_{S}^{\text{PL}}\left(\theta_{T}\right)\right)$ so that the dependence between $\theta_{S}$ and $\theta_{T}$ is explicit.
Meta Pseudo Labels (MPL) [59] is exploiting this dependence between $\theta_{S}$ and $\theta_{T}$. In other words, because the loss that the student suffers on $\mathcal{L}_{L}\left(\theta_{S}^{\text{PL}}\left(\theta_{T}\right)\right)$ is a function of $\theta_{T}$, and making a similar notational convention for the unlabeled batch $\mathcal{L}_{u}\left(\theta_{T},\theta_{S}\right)=\ell\left(\operatorname*{arg\,max}_{\xi}f_{\theta_{T}}(X_{u})|_{\xi},f_{\theta_{S}}(X_{u})\right)$, then one can further optimize $\theta_{T}$ as a function of the performance of $\theta_{S}$:
$\left\{\begin{array}[]{rl}\displaystyle\min_{\theta_{T}}&\displaystyle\mathcal{L}_{L}\left(\theta_{S}^{\text{PL}}\left(\theta_{T}\right)\right)\,,\\[3.00003pt] \text{where}&\displaystyle\theta_{S}^{\text{PL}}(\theta_{T})\in\operatorname*{arg\,min}_{\theta_{S}}\mathcal{L}_{u}\left(\theta_{T},\theta_{S}\right)\,.\end{array}\right.$ (2)
However, because the dependency between $\theta_{T}$ and $\theta_{S}$ is complicated, a practical approximation [51, 35] is obtained via $\theta_{S}^{\text{PL}}\approx\theta_{S}-\eta_{S}\cdot\nabla_{\theta_{S}}\mathcal{L}_{u}\left(\theta_{T},\theta_{S}\right)$ which leads to the practical teacher objective:
$\min_{\theta_{T}}\mathcal{L}_{L}\left(\theta_{S}-\eta_{S}\cdot\nabla_{\theta_{S}}\mathcal{L}_{u}\left(\theta_{T},\theta_{S}\right)\right)\,.$ (3)
For more information the interested reader can refer to [59].
3.3 Co-Training
Co-training relies on the existence of two different views of the instance space $\mathcal{X}$; that is, $\mathcal{X}=\mathcal{X}^{(1)}\times\mathcal{X}^{(2)}$. Hence, an instance $x\in\mathcal{X}$ has the form $x=(x^{(1)},x^{(2)})\in\mathcal{X}^{(1)}\times\mathcal{X}^{(2)}$. We say that the views $x^{(1)}$ and $x^{(2)}$ are complementary. Co-training requires two assumptions in order to work. First, it is assumed that, provided enough many examples, each separate view is enough for learning a meaningful model. The second is the conditional independence assumption [12], which states that $x^{(1)}$ and $x^{(2)}$ are conditionally independent given the label of the instance $x=(x^{(1)},x^{(2)})$. Thus, with co-training, two models $f^{(1)}$ and $f^{(2)}$ are learnt, which help each other using different information that they capture from the different views of each instance. More information is presented below.
Given a dataset $S=L\cup U$ (Section 3.1 has notation), we can have access to two different views of the dataset due to the natural partitioning of the instance space $\mathcal{X}$; that is, $S^{(1)}=L^{(1)}\cup U^{(1)}$ and $S^{(2)}=L^{(2)}\cup U^{(2)}$. Initially two models $f^{(1)}$ and $f^{(2)}$ are learnt using a supervised learning algorithm based on the labeled sets $L^{(1)}$ and $L^{(2)}$ respectively. Learning proceeds in iterations, where in each iteration a number of unlabeled instances are integrated with the labeled set by assigning pseudo-labels to them that are predicted by $f^{(1)}$ and $f^{(2)}$, so that improved models $f^{(1)}$ and $f^{(2)}$ can be learnt by retraining on the augmented labeled sets. In particular, in each iteration, each model $f^{(v)}$ predicts a label for each view $x^{(v)}\in U^{(v)}$. The top-$k$ most confident predictions by model $f^{(v)}$ are used to provide a pseudo label for the respective complementary instances and then these pseudo-labeled examples are integrated in to the labeled set for the complementary view. In the next iteration a new model is trained based on the augmented labeled set. Both views of the used instances are dropped from $U$ even if only one of the views may be used for augmenting the appropriate labeled set. This process repeats until the unlabeled sets $U^{(v)}$ are exhausted, yielding models $f^{(1)}$ and $f^{(2)}$ corresponding to the two views. Co-Training is evaluated using the joint prediction of these models which is the re-normalized element-wise product of the two predictions:
$\frac{f^{(1)}\left(x^{(1)}\right)\otimes f^{(2)}\left(x^{(2)}\right)}{\left\lVert f^{(1)}\left(x^{(1)}\right)\otimes f^{(2)}\left(x^{(2)}\right)\right\rVert_{1}}\,.$ (4)
Co-training was introduced by Blum and Mitchell [12] and their paradigmatic example was classifying web pages using two different views: (i) the bags of words based on the content that they had and (ii) the bag of words formed by the words that were used for hyperlinks pointing to these webpages. In our context the two different views are obtained via different embeddings that we obtain when we use images as inputs to different foundation models and observe the resulting embeddings.
4 Proposed Methods
4.1 View Construction
Past methods of view construction include manual feature subsetting [33], automatic feature subsetting [22, 65], random feature subsetting [15, 16], random subspace selection [16], and adversarial examples [60]. We propose a new method of view construction which is motivated by modern approaches to semi-supervised and representation learning in computer vision and natural language processing: pretext learning. Given two pretext tasks we can train a model to accomplish each task, then use their learned embedding spaces as separate views for classification. If we have chosen our pretext tasks well, then the representations will be different, compressed, and capture sufficient information to perform the desired downstream task.
While our approach is more widely applicable, it is particularly well-suited to CV. There are a plethora [38, 23, 24, 19, 49, 20, 61, 58, 42] of competitive learned representations for images. Recently it has been popular to treat these representations as task-agnostic foundation models. We show that the existence of these models eliminates the need to train new models for pretext tasks to achieve state-of-the-art performance for SSL-classification. Furthermore, it is comparatively very cheap to use the foundation models instead as not only can we forego training on a pretext task, but we are able to completely bypass the need to train a large convolutional neural network (CNN) or vision transformer (ViT). In fact, we compute the embedding for each of the images in our dataset only one time (as opposed to applying some form of data augmentation during training and making use of the embedding process during training). So, at training time the only weights we need to use at all are those in the model we train on top of the learned representation. This makes our approach incredibly lightweight compared to traditional CNN/ViT training procedures.
4.2 Meta Co-Training
In addition to our method of view construction we propose a novel algorithm for co-training which is based on Meta Pseudo Labels [59], called Meta Co-Training (MCT). Algorithm 1 presents the details of this method.
1:  Input: $L^{(1)},L^{(2)},U^{(1)},U^{(2)},f^{(1)},f^{(2)},\theta_{1},\theta_{2},T,\ell,\eta$
2:  for step $t\in\{1\dots T\}$ do
3:     // Unlabeled views must be complementary
4:     $U_{t}^{(1)}\leftarrow SampleBatch(U^{(1)})$
5:     $U_{t}^{(2)}\leftarrow getOtherView(U_{t}^{(1)},2)$
6:     // Predict soft pseudo labels
7:     $\hat{y}^{(1)}\leftarrow f^{(1)}_{\theta_{1}}(U_{t}^{(1)})$
8:     $\hat{y}^{(2)}\leftarrow f^{(2)}_{\theta_{2}}(U_{t}^{(2)})$
9:     // Sample the pseudo labels from the discrete distribution over the classes
10:     $PL^{(1)}\sim\hat{y}^{(1)}$
11:     $PL^{(2)}\sim\hat{y}^{(2)}$
12:     // MPL student update applied to both models
13:     $\theta^{\prime}_{1}\leftarrow\theta_{1}-\eta\cdot\nabla_{\theta_{1}}\ell(PL^{(2)},\hat{y}^{(1)})$
14:     $\theta^{\prime}_{2}\leftarrow\theta_{2}-\eta\cdot\nabla_{\theta_{2}}\ell(PL^{(1)},\hat{y}^{(2)})$
15:     // Sample batches from $L$ for the teacher updates
16:     $X_{t}^{(1)},Y_{t}^{(1)}\leftarrow SampleBatch(L^{(1)})$
17:     $X_{t}^{(2)},Y_{t}^{(2)}\leftarrow SampleBatch(L^{(2)})$
18:     // MPL teacher update applied to both models
19:     $\hat{y}^{\prime(1)}\leftarrow f^{(1)}_{\theta^{\prime}_{1}}(X_{t}^{(1)})$
20:     $\hat{y}^{\prime(2)}\leftarrow f^{(2)}_{\theta^{\prime}_{2}}(X_{t}^{(2)})$
21:     $h^{(1)}\leftarrow\nabla_{\theta^{\prime}_{2}}\ell(Y_{t}^{(2)},\hat{y}^{\prime(2)})^{T}\cdot\nabla_{\theta_{2}}\ell(PL^{(1)},\hat{y}^{(2)})$
22:     $h^{(2)}\leftarrow\nabla_{\theta^{\prime}_{1}}\ell(Y_{t}^{(1)},\hat{y}^{\prime(1)})^{T}\cdot\nabla_{\theta_{1}}\ell(PL^{(2)},\hat{y}^{(1)})$
23:     // Weights to be used in the next step
24:     $\theta_{1}\leftarrow\theta^{\prime}_{1}-\eta\cdot h^{(1)}\cdot\nabla_{\theta_{1}}\ell(PL^{(1)},\hat{y}^{(1)})$
25:     $\theta_{2}\leftarrow\theta^{\prime}_{2}-\eta\cdot h^{(2)}\cdot\nabla_{\theta_{2}}\ell(PL^{(2)},\hat{y}^{(2)})$
26:  end for
Algorithm 1 Meta Co-Training (Section 3.1 has notation)
At each step $t$ of the algorithm each model is first updated based on the pseudo labels the other has provided on a batch of unlabeled data (Lines 16-17). Then each model is updated to provide labels that encourage the other to predict more correctly on the labeled set based on the performance of the other (Lines 23-24). We refer the reader to Section 3.2 for a brief explanation as to why these updates make sense, or to [59] for a more complete discussion of the MPL approach. Geometrically, we can understand this as updating the parameters in the direction of increasing confidence on our pseudo labels if those labels helped the other model perform better on the labeled set, or in the opposite direction if it hurt the other model’s performance on the labeled set. Note that contrary to traditional co-training, no new pseudo-labeled data are added to the labeled set, neither are any data removed from the unlabeled set after each step.
5 Experiments
Towards evaluating the proposed methods of view construction and co-training we first constructed views of the data utilizing five different learned representations. Our primary comparisons were on the widely-used ImageNet [31] classification task; we provide experimental results on additional image classification tasks. To produce the views used to train the classifiers during co-training and meta co-training we used the embedding space of five representation learning architectures: The Masked Autoencoder (MAE) [42], DINOv2 [58], SwAV [19], EsViT [49], and CLIP [61]. We selected the models which produce the views as they have been learned in an unsupervised way, have been made available by the authors of their respective papers for use in PyTorch, and have been shown to produce representations that are appropriate for ImageNet classification.
On the Sufficiency of the Views.   Motivated by the analysis and experiments of [34] we made an effort to verify the sufficiency and independence of the representations learned by the aforementioned models. Sufficiency is fairly easy to verify. If we choose a reasonable sufficiency threshold for the task, for ImageNet say $75\%$ top-1 accuracy, then we can train simple models on the views of the dataset we have constructed and provide examples of functions which demonstrate the satisfaction of the property. We tested a single linear layer with softmax output (Table 1), and a simple 3-layer multi-layer perceptron (MLP) with 1024 neurons per layer and dense skip connections (previous layer outputs were concatenated to input for subsequent layers) (Table 2) to provide a lower bound on the sufficiency of the views on the standard subsets of the ImageNet labels for semi-supervised classification.
Model 1% 10% 100%
MAE 1.9 3.2 73.5
DINOv2 78.1 82.9 86.3
SwAV 12.1 41.1 77.9
EsViT 69.1 74.4 81.3
CLIP 74.1 80.9 85.4
Table 1: Linear probe evaluation for views. Top-1 accuracy on different subsets of the ImageNet data are shown.
Model 1% 10%
MAE 23.4 48.5
DINOv2 78.4 82.7
SwAV 12.5 32.6
EsViT 71.3 75.8
CLIP 75.2 80.9
Table 2: MLP performance on individual views. Top-1 accuracy on different subsets of the ImageNet data are shown.
From these experiments we can see that if we were to pick the $75\%$ top-1 accuracy threshold, which we believe indicates considerable skill without presenting too high of a bar, then CLIP and DINOv2 provide sufficient views for both the $1\%$ and $10\%$ subsets.
On the Independence of the Views.   To test the independence of the views generated by the representation learning models, we trained an identical MLP architecture to predict each view from each other view; Table 3 presents our findings, explained below.
MAE DINOv2 SwAV EsViT CLIP
MAE - 0.139 0.142 0.137 0.129
DINOv2 0.110 - 0.132 0.135 0.130
SwAV 0.116 0.147 - 0.137 0.126
EsViT 0.112 0.140 0.116 - 0.135
CLIP 0.110 0.146 0.136 0.156 -
Table 3: Pairwise translation performance of linear probe on the ImageNet dataset. A linear classifier is trained on the output of an MLP which is trained to predict one view (columns) from another view (rows) by minimizing MSE. The top-1 accuracy (%) of the linear classifier is reported. Both the MLP and the linear classifier have access to the entire embedded ImageNet training set.
Given as input the view to predict from (rows of Table 3) the MLPs were trained to reduce the mean squared error (MSE) between their output and the view to be predicted (columns of Table 3). We then trained a linear classifier on the outputs generated by the MLP given every training set embedding from each view. Had the model faithfully reconstructed the output view given only the input view then we would expect the linear classifier to perform similarly to the last column of Table 1. The linear classifiers never did much better than a random guess on the ImageNet class achieving at most 0.156% accuracy. Thus, while we cannot immediately conclude that the views are independent, it is clearly not trivial to predict one view given any other. We believe that this is compelling evidence for the independence of different representations.
Co-Training Experiments on ImageNet.   Having constructed at least two strong views of the data, and suspecting that these views are independent we hypothesize that co-training will work on this data. In Table 3(a) we show the results of applying the standard co-training algorithm [12]. At each iteration of co-training the model made a prediction for all of the instances in $U$, following (4). We assigned labels to the 10% of the total original unlabeled data with most confident predictions for both models. When confident predictions conflicted on examples, we returned those instances to the unlabeled set, otherwise they entered the labeled set of the other model with the assigned pseudo-label.
Comparing co-training accuracy on ImageNet 10% (resp. 1%) shown in Table 3(a) compared to MLP accuracy shown in Table 2, we observe that co-training top-1 accuracy is on average 18.7% (resp. 16.8%) higher than MLP accuracy. Moreover, co-training top-1 accuracy is 2.4% (resp. 1.7%) higher overall. In other words, co-training resulted in a significant increase in prediction accuracy but not after any iterations of co-training have occurred; see Figures 1 and 2. Furthermore, predicting according to (4) yielded better accuracy than the predictions of each individual model. In addition, in our experiments on ImageNet, as co-training iterations proceeded top-1 accuracy decreased. Later we show that this was not always the case (see Figure 9 in the appendix). The decrease was less pronounced when the views perform at a similar level (see Figure 3). Despite the poor performance of the pseudo-labeling during co-training, just computing the re-normalized product of the two best performing views performed as well as the previous SotA method for semi-supervised classification on the ImageNet-10% task (see Co-Training in Table 5). The performance of co-training as a method was rather disappointing in that the algorithm is unable to leverage pseudo-labels to improve the accuracy on this task; however, the results do show that our view-construction method has merit. Given the difficulty of translating between the views, and the fact that the accuracy yielded by (4) is greater than its constituent views, we believe the poor performance of the pseudo-labeling step in co-training was due to the insufficiency of at least one view rather than view-dependence.
Figure 1: Top-1 accuracy of co-training iterations for the ImageNet 10% dataset. The average over all 10 pairs, the best pair, and the worst pair at each iteration are shown.
Figure 2: Top-1 accuracy of co-training iterations for the ImageNet 1% dataset. The average over all 10 pairs, the best pair, and the worst pair at each iteration are shown.
Figure 3: Top-1 accuracy of co-training iterations on the CLIP and DINOv2 views for the ImageNet 10% dataset.
Figure 4: Top-1 accuracy of co-training iterations on the CLIP and DINOv2 views for the ImageNet 1% dataset.
DINOv2 SwAV EsViT CLIP
MAE 81.8 (75.5) 30.5 (11.4) 77.1 (66.2) 78.8 (66.8)
DINOv2 78.5 (74.5) 83.3 (78.9) 85.1 (80.1)
SwAV 70.6 ( 64.9) 75.5 (67.4)
EsViT 82.3 ( 76.9)
(a) Co-Training
DINOv2 SwAV EsViT CLIP
MAE 81.2 (73.8) 37.6 (8.1) 73.6 ( 63.8) 78.6 (63.5)
DINOv2 77.3 ( 70.7) 83.4 (79.2) 85.2 (80.7)
SwAV 74.4 (67.3) 77.1 (67.4)
EsViT 82.4 (77.5)
(b) Meta Co-Training
Table 4: Co-training and meta co-training top-1 accuracy of view combinations on 10% (1%) of ImageNet labels. As co-training does not depend on the order of the views, we show only upper-diagonal entries of the pairwise comparison. Diagonal entries would correspond to views which are trivially dependent. Note that all $\binom{5}{2}=10$ different combinations of views are shown in each table.
Reference Model Method ImageNet-1% ImageNet-10%
[49] EsViT (Swin-B, W=14) Linear 69.1 74.4
[42] MAE (ViT-L) MLP 23.4 48.5
[61] CLIP (ViT-L) Fine-tuned 80.5 84.7
[58] DINOv2 (ViT-L) Linear 78.1 82.9
[24] SimCLRv2 (ResNet152-w2) Fine-tuned 74.2 79.4
[19] SwAV (RN50-w4) Fine-tuned 53.9 70.2
[60] Deep Co-Training (ResNet-18) Co-Training - 53.5
[17] Semi-ViT (ViT-L) Self-trained 77.3 83.3
[52] REACT (ViT-L) REACT 81.6 85.1
[59] MPL (EfficientNet-B6-Wide) MPL - 73.9
[12] Co-Training (MLP)* Co-Training 80.1 85.1
MCT (MLP)* - ours MCT 80.7 85.8
Table 5: Performance of different approaches on ImageNet dataset. An asterisk (*) indicates models that were trained on top of embeddings generated by much larger models. During training we do not need to alter the parameters of these larger models, and these larger models need not see data used for downstream classification during their training, however training of downstream models would not have been possible without the embeddings. Models on the lower half of the table use unlabeled data during classification training.
Meta Co-Training Experiments on ImageNet.   To draw fair comparisons we fix the model architecture and view set from our experiments on co-training. As in [59] a supervised loss is optimized jointly with the MPL loss. We found that beginning with a warmup period in which the model trained in a strictly supervised way expedited training. Unlike MPL, we did not use Unsupervised Data Augmentation [75], because we embedded the dataset only once, and due to the nature of the pretext tasks that were used to generate the views we did not believe data augmentation would have much of an effect. Details of our training recipe appear in Table 10.
Comparing meta co-training accuracy on ImageNet 10% (resp. 1%) shown in Table 3(b) compared to MLP accuracy shown in Table 3(a), we observe that meta co-training top-1 accuracy yielded a difference on average 0.7% (resp. -0.6%) compared to co-training top-1 accuracy. Moreover, meta co-training top-1 accuracy is 0.1% (resp. 0.6%) higher overall. Note that on all view pairs in which co-training beats meta co-training, with the exception of the $10\%$ (EsViT, MAE) pair, one of the views alone was better than co-training. The single DINOv2 view outperforms all such view pairs. These are all cases in which one of the views was significantly weaker than the other. For the experiments in which we had two strong (e.g., not SwAV or MAE) views, meta co-training always had better performance than co-training. Finally, we take the four views which had the greatest performance and constructed two views out of them by concatenating them together. We constructed one view CLIP |||| EsViT and the second DINOv2 |||| SwAV and measured the performance of co-training and meta co-training on these views in Table 6. Interestingly, co-training does not benefit from these larger views; see Table 6.
Method 1% 10%
Co-Training 80.0 84.8
MCT 80.5 85.8
Table 6: Meta co-training and co-training evaluated on the ImageNet dataset using the CLIP |||| EsViT and DINOv2 |||| SwAV views.
The additional information may have allowed the individual models to overfit their training data quickly. During meta co-training we observed that after one model reaches 100% training accuracy, its validation accuracy can still improve by learning to label for the other model. Possibly due to the rapid overfitting due to the increased view size (and consequently increased parameter count) the 1% split did not show improvement for either model.
Dataset
Method Flowers102 Food101 FGVCAircraft
REACT (ViT-L) 97.0 85.6 57.1
Co-Training (MLP) 99.2 94.7 36.4
MCT (MLP) 99.6 94.8 40.1
Table 7: Additional Comparisons to REACT [52]. The authors include experiments for zero-shot performance and 10% of available labels. Results above are for 10% of labels. For Flowers102 this is 1-shot performance.
Dataset
Method iNaturalist Food101
Semi-ViT (ViT-B) 67.7 (32.3) 84.5 (60.9)
Co-Training (MLP) 59.7 (29.5) 94.7 (83.9)
MCT (MLP) 76.0 (58.1) 94.8 (91.7)
Table 8: Additional Comparisons to Semi-ViT using 10% (1%) of the available labels for training. For the iNaturalist task we use only the 1010 most frequent classes as in [17].
Additional Datasets.   In Tables 7 and 8 we compare meta co-training and co-training to existing SotA approaches that leverage unlabeled data directly during training (e.g., not just part of a pretext task). These experiments support our hypothesis that the primary cause of the failure of co-training is the insufficiency of one or more views. The closer to equivalent in performance individual views are, the less the performance suffers. When the two views perform similarly and well, the performance improves. Section A in the appendix provides more information on these additional datasets; Figure 9 provides an example where co-training performs as expected. Meta co-training is in general more robust to view imbalance and provides better results.
6 Conclusion and Future Work
We presented a novel method of view construction for co-training approaches based on pretext tasks. We showed that the views themselves likely contain independent information and that models trained on these views can complement each other to produce stronger classifiers. However, these views appear to be insufficient for co-training on the ImageNet task. We proposed meta co-training, a novel algorithm for co-training to address the challenges faced by the co-training algorithm. We showed that our algorithm outperforms co-training in general, and establishes a new SotA top-1 classification accuracy on the ImageNet-10% dataset as well as new SotA few-shot top-1 classification accuracy on several other fine-grained image classification datasets.
For future work we expect that fine-tuning the representation learners that we used for co-training and meta co-training would yield better results; which should be in line to observations from papers proposing foundation models that report performance benefits of fine-tuning those models on downstream tasks. Furthermore, we did not make use of the large available unlabeled datasets for computer vision in this study. We suspect that giving our models access to a large repository of unlabeled data as well as the entirety of the labeled data for any dataset on which we performed experiments would result in a performance boost. Finally, our list of pretext tasks was not exhaustive. It is an interesting question for future work which other tasks and models might yield useful views.
Acknowledgements.   This material is based upon work supported by the National Science Foundation under Grant No. ICER-2019758. This work is part of the NSF AI Institute for Research on Trustworthy AI in Weather, Climate, and Coastal Oceanography (AI2ES).
Some of the computing for this project was performed at the OU Supercomputing Center for Education & Research (OSCER) at the University of Oklahoma (OU).
Appendix A Further Analysis of Meta Co-Training and Co-Training
Figures 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, and 15 include the top-1 accuracy of co-training using the best (CLIP and DINO) views found for the ImageNet task. In all cases except Figure 9, co-training reaches its maximum performance in the first iteration. Note that there is no 1% run for Flowers102 because there are only 10 examples for each class in the training set (see the discussion in Section B).
A.1 Further Analysis of Meta Co-Training Experiments on ImageNet
In Figures 5 and 6 are shown aggregate results comparable to those in Figures 1 and 2 in the main text, but for meta co-training. The min, max, and mean performance of any of the ten pairs is shown. Note that these are not three runs, but aggregate statistics over all runs, so the minimum and maximum points on the graph for one step may not correspond to the same pair as another step.
Figure 5: Aggregate statistics of meta co-training across all views. Models are trained on 10% of the ImageNet labels. The maximum performance of any pair, the minimum performance of any pair, and the average of all 10 pairs is shown.
Figure 6: Aggregate statistics of meta co-training across all views. Models are trained on 1% of the ImageNet labels. The maximum performance of any pair, the minimum performance of any pair, and the average of all 10 pairs is shown.
In Figures 7 and 8 are shown results comparable to those in Figures 3 and  4 in the main text, but for meta co-training. Unlike co-training, meta co-training performance increases after the supervised warmup.
Figure 7: Meta co-training using the CLIP and DINOv2 views as a function of the training step. Models are trained on 10% of the ImageNet labels.
Figure 8: Meta co-training using the CLIP and DINOv2 views as a function of the training step. Models are trained on 1% of the ImageNet labels.
A.2 Co-Training and Meta Co-Training on Food101
In Figures 9 and  10 we show the top-1 accuracy of the iterations of co-training on the Food101 classification task. Co-training performs as expected with the pseudo labels allowing the other model to improve in the 10% case, but in the 1% case it appears that one or more of the views was insufficient and the performance decreases rapidly after the first iteration. Co-training achieves top-1 accuracy of 94.7% (resp., 83.9%) when 10% (resp., 1%) of the available labels for training are used from the Food101 dataset.
Figure 9: The top-1 accuracy of the co-training predictions for each iteration of co-training. 10% of available Food101 labels are used for training. The method exhibits performance improvement over multiple iterations of pseudo-labeling and retraining.
Figure 10: The top-1 accuracy of the co-training predictions for each iteration of co-training. 1% of available Food101 labels are used for training.
Similarly, meta co-training achieves top-1 accuracy of 94.8% (resp., 91.7%) when 10% (resp., 1%) of the available training examples of the Food101 dataset are used.
These results are compared against each other as well as against REACT and Semi-ViT in Tables 7 and 8 respectively.
A.3 Co-Training and Meta Co-Training on iNaturalist
In Figures 11 and 12 we show the top-1 accuracy of the iterations of co-training on the iNaturalist dataset. In both splits we see a steady decline in top-1 accuracy after the first iteration. Co-training achieves top-1 accuracy of 59.7% (resp., 29.5%) when only 10% (resp., 1%) of the training examples of the iNaturalist dataset are used.
Figure 11: The top-1 accuracy of the co-training predictions for each iteration of co-training. Only the 1010 most frequent classes are included. 10% of available iNaturalist labels from those classes are used for training.
Figure 12: The top-1 accuracy of the co-training predictions for each iteration of co-training. Only the 1010 most frequent classes are included. 1% of available iNaturalist labels from those classes are used for training.
Along these lines meta co-training achieves 76.0% (resp., 58.1%) when only 10% (resp., 1%) of the training examples of the iNaturalist dataset are used.
These results are compared against each other as well as against REACT and Semi-ViT in Tables 7 and 8 respectively.
A.4 Co-Training and Meta Co-Training on CTFGVCAircraft
In Figures 13 and 14 we show the top-1 accuracy of the iterations of co-training on the CTFGVCAircraft dataset. In both cases both views have low accuracy, so it is unsurprising that in subsequent iterations the accuracy decreases rapidly. In particular, co-training achieves top-1 accuracy of 36.4% (resp., 21.6%) when only 10% (resp., 1%) of the training examples of the CTFGVCAircraft dataset are used.
Figure 13: The top-1 accuracy of the co-training predictions for each iteration of co-training. 10% of available FGVCAircraft labels are used for training.
Figure 14: The top-1 accuracy of the co-training predictions for each iteration of co-training. 1% of available FGVCAircraft labels are used for training.
Along these lines meta co-training achieves 40.1% (resp., 21.5%) when only 10% (resp., 1%) of the training examples of the CTFGVCAircraft dataset are used.
The results corresponding to the 10% case of co-training and meta co-training are compared against REACT in Table 7, since the authors of [52] report results on zero-shot and 10% cases only. However, for completeness, we have decided to include Figure 14 which corresponds to the experiments that we did using only 1% of the labeled training data available for the CTFGVCAircraft dataset.
Dataset #train #test #classes
FGVCAircraft $3$ $3$ 100
Flowers102 1020 6149 102
Food101 75,750 25,250 101
iNaturalist 175,489 29,083 1010
ImageNet 1,281,167 50,000 1000
Table 9: Sizes of datasets used. Subsets drawn for training are always drawn in a stratified way to maintain the class distribution of the original data. For ImageNet we use the split published with the SimCLRv2 [24] repository.
config MCT Co-Training
optimizer Adam Adam
optimizer momentum $\beta_{1},\beta_{2}=0.9,0.999$ $\beta_{1},\beta_{2}=0.9,0.999$
base learning rate 1e-4 1e-4
batch size 4096 4096
learning rate schedule Reduce On Plateau Reduce On Plateau
schedule patience 10 steps 10 steps
decay factor 0.5 0.5
training steps 1000 200 (per iteration)
warmup steps 200 n/a
Table 10: Hyperparmeter configuration for meta co-training and co-training. Other models used (MLP, linear) are trained for the same number of steps as meta co-training, with the same optimization parameters.
A.5 Co-Training and Meta Co-Training on Flowers102
As explained earlier, the Flowers102 dataset has only 10 training examples for each class and hence 10% corresponds to 1-shot learning. In this context, in Figure 15 we give the performance of the iterations of co-training on the Flowers102 dataset. Even though the DINO view performs very well in the first iteration, there is a large gap between the performance of the CLIP view and the DINO view which leads to decreased performance in subsequent iterations. Co-training achieves top-1 accuracy of 99.2% when only 10% of the training examples of the Flowers102 dataset are used for training (i.e., a single labeled image per class).
In the same context of 1-shot learning, meta co-training achieves top-1 accuracy of 99.6%.
Both of the above results are compared against each other and against REACT in Table 7.
Figure 15: The top-1 accuracy of the co-training predictions for each iteration of co-training. 10% of available Flowers102 labels are used for training.
Appendix B Further Information on the Datasets and the Splits
Table 9 includes a description of the different sizes of the datasets used for training and testing, as well as the number of cases for each dataset. Note that in the 10% experiments on the Flowers102 dataset each class has only 1 label. All datasets are approximately class-balanced with the exception of the iNaturalist dataset. In all cases we maintain the original class distribution when sampling subsets. All subsets are created with seeded randomness of a common seed (13) which ensures a fair comparison between methods.
Appendix C Hyperparameter Configuration
The relevant hyperparameters for training are included in Table 10 for co-training and meta co-training (MCT).

### v5

@@
 Abstract
-In many practical computer vision scenarios unlabeled data is plentiful, but labels are scarce and difficult to obtain.
+In many critical computer vision scenarios unlabeled data is plentiful, but labels are scarce and difficult to obtain.
 As a result, semi-supervised learning which leverages unlabeled data to boost the performance of supervised classifiers have received significant attention in recent literature.
-One major class of semi-supervised algorithms is co-training.
-In co-training two different models leverage different independent and sufficient “views” of the data to jointly make better predictions.
-During co-training each model creates pseudo labels on unlabeled points which are used to improve the other model.
-We show that in the common case when independent views are not available we can construct such views inexpensively using pre-trained models.
-Co-training on the constructed views yields a performance improvement over any of the individual views we construct and performance comparable with recent approaches in semi-supervised learning, but has some undesirable properties.
-To alleviate the issues present with co-training we present Meta Co-Training which is an extension of the successful Meta Pseudo Labels approach to two views.
-Our method achieves new state-of-the-art performance on ImageNet-10% with very few training resources, as well as outperforming prior semi-supervised work on several other fine-grained image classification datasets.
+One representative class of semi-supervised algorithms are co-training algorithms.
+Co-training algorithms leverage two different models which have access to different independent and sufficient representations or “views” of the data to jointly make better predictions.
+Each of these models creates pseudo-labels on unlabeled points which are used to improve the other model.
+We show that in the common case where independent views are not available, we can construct such views inexpensively using pre-trained models.
+Co-training on the constructed views yields a performance improvement over any of the individual views we construct and performance comparable with recent approaches in semi-supervised learning.
+We present Meta Co-Training, a novel semi-supervised learning algorithm, which has two advantages over co-training: (i) learning is more robust when there is large discrepancy between the information content of the different views, and (ii) does not require retraining from scratch on each iteration.
+Our method achieves new state-of-the-art performance on ImageNet-10% achieving a $\sim 4.7\%$ reduction in error rate over prior work.
+Our method also outperforms prior semi-supervised work on several other fine-grained image classification datasets.
 1 Introduction
-In many practical machine learning scenarios we have access to a large amount of unlabeled data, and relatively few labeled data points for the task on which we would like to train.
-For generic computer vision (CV) tasks there are large well-known and open source datasets with hundreds of millions or even billions of unlabeled images [62, 63].
-In contrast, labeled data is usually scarce and expensive to obtain requiring many human-hours to generate.
-Hence, in this context, semi-supervised learning (SSL) methods are becoming increasingly popular as they rely on training useful learning models by using small amounts of labeled data and large amounts of unlabeled data.
+In many critical machine learning scenarios we have access to a large amount of unlabeled data, and relatively few labeled data points for a given task.
+For standard computer vision (CV) tasks there are large well-known and open source datasets with hundreds of millions or even billions of unlabeled images [47, 46].
+In contrast, labeled data is usually an order of magnitude more scarce and otherwise expensive to obtain requiring many human-hours to generate.
+In this context, semi-supervised learning (SSL) methods are useful as they rely on training more performant models, using small amounts of labeled data and large amounts of unlabeled data.
 Not unrelated to, and not to be confused with the idea of semi-supervised learning, is self-supervised learning.
@@
 In self-supervised learning a model is trained with an objective that does not require a label that is not evident from the data itself.
-Self-supervised learning was popularized by the BERT [32] model for generating word embeddings for natural language processing (NLP) tasks.
+Self-supervised learning was popularized by the BERT model [18] for generating word embeddings for natural language processing (NLP) tasks.
 BERT generates its embeddings by solving the pretext task of masked word prediction.
-Pretext tasks present unsupervised objectives that are not directly related to any particular supervised objective we might desire to solve, but rather are solved in the hope of learning a suitable representation to more easily solve a downstream task.
-Due to the foundational nature of these models to the success of many models in NLP, or perhaps the fact that they serve as a foundation for other models to be built upon, these pretext learners are often referred to as foundation models.
-Inspired by foundation models from NLP, several pretext tasks have been proposed for CV with associated foundation models that have been trained to solve them [8, 38, 20, 42, 23, 49, 19, 58, 24, 43].
-Unlike NLP, the learned representations for images are often much smaller than the images themselves which yields the additional benefit of reduced computational effort when using the learned representations.
-Typically, SSL requires the learner to make predictions on a large amount of unlabeled data as well as the labeled data during training.
-Furthermore, SSL often requires that the learner be re-trained [12, 76, 50, 72, 48, 55, 39, 70, 74, 40, 17] partially, or from scratch.
-Given the computational benefits of learning from a frozen representation, combining SSL and self-supervised learning is a natural idea.
-Additionally, the representation learner allows us to compress our data to a smaller size.
-This compression is beneficial not only because of the reduction in operations required that is associated with a smaller model, but also because it allows us to occupy less space on disk, in RAM, and in high bandwidth memory on hardware accelerators.
-These benefits are particularly pronounced for co-training [12] in which two different “views” of the data must be obtained or constructed and then two different classifiers must be trained and re-trained iteratively.
-Seldom does a standard problem in machine learning present two independent and sufficient views of the problem to be leveraged for co-training.
-As a result, if one hopes to apply co-training one usually needs to construct two views from the view that they have.
-To take advantage of the computational advantages of representation learning, and the large body of self-supervised literature, we propose using the embeddings produced by foundation models trained on different pretext tasks as the views for co-training.
-To our knowledge we are the first to propose using frozen learned feature representations as views for co-training classification, though we think the idea is very natural.
-We show that views constructed in this way appear to satisfy the necessary conditions for the effectiveness of co-training.
-Unfortunately, while we can boost performance by training different models on two views, co-training performs sub-optimally during its pseudo-labeling step.
-To more effectively leverage two views we propose a novel method for co-training based on Meta Pseudo Labels [59] called Meta Co-Training in which each model is trained to provide better pseudo labels to the other given only its view of the labeled data and the performance of the other model on the labeled set.
-We show that our approach provides an improvement over both co-training and the current state-of-the-art (SotA) on the ImageNet-10% dataset, as well as it establishes new SotA few-shot performance on several other fine-grained image classification tasks.
-Summary of Contributions.   Our contributions are therefore summarized as follows.
+Pretext tasks present unsupervised objectives that are not directly related to any particular supervised objective we might want to solve, but rather are solved in the hope of learning a suitable representation to more easily solve a downstream task.
+These pretext learners are often referred to as foundation models.
+Several pretext tasks have been proposed for CV to train associated foundation models that solve them [26, 12, 34, 10, 42, 27].
+The learned representations for images are often much smaller than the images themselves.
+As a consequence, this yields the additional benefit of reduced computational cost when using the learned representations.
+SSL algorithms [33, 5, 44, 52, 43, 49, 51, 4] involve generating pseudo-labels to serve as weak supervision on unlabeled data and then learning from those pseudo-labels.
+This weak supervision can either be used to perform consistency regularization, or entropy minimization, or both.
+Consistency regularization is to enforce that different augmented versions of the same instance are assigned the same label.
+Entropy minimization is to enforce that all instances are assigned a label with high confidence.
+Typically consistency regularization is achieved by minimizing a consistency loss between pairs of examples that ought to have the same label [51, 4, 49, 3], and entropy minimization is achieved by iterative re-training [33, 5, 44] or sharpening labels [43, 4, 49, 3].
+One particular class of SSL algorithms are co-training [5] algorithms in which two different “views” of the data must be obtained or constructed, and then two different models must be trained and re-trained iteratively.
+These two views yield models which capture different patterns in their input and thus, informally speaking, are more likely to fail independently.
+This independence of failure is leveraged so that each model can provide useful labels to the other.
+Standard benchmark problems in machine learning do not present two views of the problem to be leveraged for co-training.
+If one hopes to apply co-training, one needs to construct two views from the single view that they have access to.
+We show that constructing views which satisfy the conditions necessary for co-training is relatively simple.
+Training different models on the two constructed views can boost performance.
+Unfortunately, the classical co-training algorithm (CT) fails to utilize pseudo-labels effectively on benchmark tasks.
+We propose a novel SSL method which more effectively leverages pseudo-labels.
+We call our algorithm Meta Co-Training (MCT), because it leverages two views to produce good pseudo-labels as part of a bi-level optimization.
+In MCT each model is trained to provide better pseudo-labels to the other model, given only its view of the data and the performance of the other model on a labeled set.
+Similarly to CT, our approach utilizes multiple views to train models which are diverse.
+These models have an advantage when teaching (and learning from) each-other because they will learn different patterns from their differing input data.
+We show that our approach provides an improvement over both CT and the current state-of-the-art (SoTA) method [37] on the ImageNet-10% dataset, as well as it establishes new SoTA few-shot performance on several other fine-grained image classification tasks.
+Summary of Contributions.
+Our contributions are summarized as follows.
   1. 1.
-    Use of embeddings obtained from different pre-trained models as different views for co-training.
+    We propose Meta Co-Training, an SSL algorithm which appears to be more robust than co-training and does not require iterative re-training from initialization.
   2. 2.
-    We boost the performance of standard co-training by proposing a method that relies on Meta Pseudo Labels [59], called Meta Co-Training.
+    We establish SoTA top-1 ImageNet-10% [17] accuracy and other few-shot benchmark classification tasks.
   3. 3.
-    We establish new state-of-the-art top-1 accuracy on ImageNet-10% [31] and new state-of-the-art top-1 few-shot accuracy on several other CV classification tasks.
-  4. 4.
-    We make our implementation publicly available:
-    https://github.com/JayRothenberger/Meta-Co-Training
+    We make our implementation publicly available for others to use and build upon. https://github.com/JayRothenberger/Meta-Co-Training
 Outline of the Paper.
-In Section 2 we discuss related work.
-In Section 3 we establish notation and provide more information on co-training and meta pseudo labels.
-In Section 4 we describe our proposed methods of view construction and of meta co-training.
-In Section 5 we present and discuss our experimental findings on ImageNet [31], Flowers102 [57], Food101 [14], FGVCAircraft [54], iNaturalist [44], and iNaturalist 2021 [45].
-In Section 6 we conclude with a summary and ideas for future work.
-2 Related Work
-While SSL [21, 78, 67, 73] has become very relevant in recent years with the influx of the data collected by various sensors, the paradigm and its usefulness have been studied at least since the 1960’s; e.g., [41, 29, 64, 36].
-Among the main approaches for SSL that are not directly related to our work, we find techniques that try to identify the maximum likelihood parameters for mixture models [30, 7], constrained clustering [9, 28, 69, 68, 18, 3], graph-based methods [11, 10, 53, 6], as well as purely theoretical probably approximately correct (PAC)-style results [4, 37, 25].
-Below we discuss SSL approaches closer to our work.
-Self-Training and Student-Teacher Frameworks.
-One of the earliest ideas for SSL is that of self-training [64, 36].
-This idea is still very popular with different student-teacher frameworks [79, 76, 59, 50, 72, 66, 55, 39, 17].
-This technique is also frequently referred to as pseudo-labeling [48] as the student tries to improve its performance by making predictions on the unlabeled data, then integrating the most confident predictions to the labeled set, and subsequently retraining using the broader labeled set.
-In addition, one can combine ideas; e.g., self-training with clustering [46], or create ensembles via self-training [47], potentially using graph-based information [53].
-An issue with these approaches is that of confirmation bias [2], where the student cannot improve a lot due to incorrect labeling that occurs either because of self-training, or due to a fallible teacher.
-Co-Training.
-Another paradigmatic family of SSL algorithms are co-training [12] algorithms, where instances have two different views for the same underlying phenomenon.
-Co-training has been very successful with wide applicability [34] and can outperform learning algorithms that use only single-view data [56].
-Furthermore, the generalization ability of co-training is well-motivated theoretically [12, 27, 5, 26].
-There are several natural extensions of the base idea, such as methods for constructing different views [71, 15, 22, 65, 16], including connections to active learning [33], methods that are specific to deep learning [60], and other natural extensions that exploit more than two views on data; e.g., [77].
+In Section 2 we establish notation and provide background information necessary to understand our method.
+In Section 3 we describe our proposed method, meta co-training.
+In Section 4 we present our experimental findings on ImageNet [17] and discuss how our method compares to relevant baseline approaches.
+We perform additional experiments on Flowers102 [41], Food101 [6], FGVCAircraft [39], iNaturalist [29], and iNaturalist 2021 [28] datasets.
+Additional details such as hyperparameters are provided in Appendix B.
+In Section 5 we discuss related work with emphasis on semi-supervised methods, including co-training methods.
+In Section 6 we conclude with a summary and directions for future work.
 2 Background
-We briefly describe two important components of semi-supervised learning that are used jointly in order to devise our proposed new technique.
-However, before we do so, we also define basic notions so that we can eliminate ambiguity.
-3.1 Notation
-Supervised learning is based on the idea of learning a function, called a model, $f$ belonging to some model space $\mathcal{F}$, using training examples $(x,y)$ that exemplify some underlying phenomenon.
-Training examples are labeled instances; i.e., $x\in\mathcal{X}$ is an instance drawn from some instance space $\mathcal{X}$ and $y$ is a label from some label set $\mathcal{Y}$.
-This relationship between $\mathcal{X}$ and $\mathcal{Y}$ can be modeled with a probability distribution $D$ that governs $\mathcal{X}\times\mathcal{Y}$.
-We write $(x,y)\sim D$ to indicate that a training example $(x,y)$ is drawn from $\mathcal{X}\times\mathcal{Y}$ according to $D$.
-We use $\mathbf{1}\left\{\mathcal{A}\right\}$ as an indicator function of the event $\mathcal{A}$; that is, $\mathbf{1}\left\{\mathcal{A}\right\}$ is equal to 1 when $\mathcal{A}$ holds, otherwise 0.
-For vectors $\alpha,\beta\in\mathbb{R}^{d}$ we indicate their Hadamard (element-wise) product using $\otimes$; that is, $\gamma=\alpha\otimes\beta$, where $\gamma_{j}=\alpha_{j}\cdot\beta_{j}$ for $j\in\{1,\ldots,d\}$.
-As is standard in classification problems, we let model $f$ be a function of the form $f\colon\mathcal{X}\rightarrow\Delta^{|\mathcal{Y}|}$ where $\Delta^{|\mathcal{Y}|}$ is the $|\mathcal{Y}|$-dimensional unit simplex.
-Equivalently $f$ can be understood as a probability distribution $\mathbf{Pr}(y\ |\ x)$.
-We use $f(x)|_{j}$ to denote the $j$-th value of $f$ in the output in such a case.
-It is common for $\mathcal{F}$ to be a parameterized family of functions (e.g.
-artificial neural networks).
-We use $\theta$ as a vector that holds such parameters, and we denote the function parameterized by $\theta$ as $f_{\theta}$.
-An individual prediction $f_{\theta}(x)$ for an instance $x$ is evaluated via the use of a loss function $\ell$ that determines the quality of the prediction $f_{\theta}(x)$ with respect to the actual value $y$ that should have been predicted.
-For classification problems a natural way of defining loss functions is of the form $\ell\colon\mathcal{Y}\times\Delta^{|\mathcal{Y}|}\rightarrow\mathbb{R}_{+}$.
+Notation.
+The models that we learn are functions of the form $f\colon\mathcal{X}\rightarrow\Delta^{|\mathcal{Y}|}$ where $\Delta^{|\mathcal{Y}|}$ is the $|\mathcal{Y}|$-dimensional unit simplex.
+We use $f(x)|_{j}$ to denote the $j$-th value of $f$ in the output.
+When we want to explicitly refer to a network $f$ parameterized by $\theta$ (e.g., in deep neural networks), we write $f_{\theta}$.
 For the entirety of this text $\ell$ will refer to the cross-entropy loss which is defined as $\ell\left(y,f_{\theta}(x)\right)=-\sum_{\xi\in\mathcal{Y}}\mathbf{1}\left\{\xi=y\right\}\log\left(f_{\theta}(x)|_{\xi}\right)$.
-We measure the fitness of a model $f_{\theta}$ by its (true) risk $R_{D}\left(f_{\theta}\right)=\mathbf{E}_{(x,y)\sim D}\left[\ell\left(y,f_{\theta}(x)\right)\right]$.
-While we want to create a model $f_{\theta}$ that has low true risk, in principle we apply empirical risk minimization (ERM) which yields a model $f_{\theta}$ that minimizes the risk over a set of training examples.
-That is, we use empirical risk as a proxy for the true risk.
-Given a set $S=\left\{(x_{i},y_{i})\right\}_{i=1}^{\left\lvert S\right\rvert}$, we denote the empirical risk of $f_{\theta}$ on $S$ as $\widehat{R}_{S}\left(f_{\theta}\right)=\frac{1}{\left\lvert S\right\rvert}\sum_{i=1}^{\left\lvert S\right\rvert}\ell\left(y,f_{\theta}(x)\right)$.
-To ease notation, when the inputs to $\ell$ are a batch of examples $(X,Y)$, we denote $\widehat{R}_{(X,Y)}\left(f_{\theta}\right)$ as simply $\ell(Y,f_{\theta}(X))$.
-Similarly, when we write the arg max of a batch of vectors we refer to the batch of the arg max of the individual values.
-In general, ERM can be computationally hard (e.g., [13, 1]); hence, we often rely on gradient-based optimization techniques, such as stochastic gradient descent (SGD), that can take us to a local optimum efficiently.
-We use $\eta$ as the learning rate and $T$ as the maximum number of steps (or updates) in such a gradient-based optimization procedure.
+We use $\eta$ as the learning rate and $T$ as the maximum number of steps (or updates) in a gradient-based optimization procedure.
+Finally, we use $\mathbf{1}\left\{\mathcal{A}\right\}$ as an indicator function of an event $\mathcal{A}$; that is, $\mathbf{1}\left\{\mathcal{A}\right\}$ is equal to 1 when $\mathcal{A}$ holds, otherwise 0.
 Semi-Supervised Learning.
 We are interested in situations where we have a large pool of unlabeled instances $U$ as well as a small portion $L$ of them that is labeled.
-In other words, we assume that the learning algorithm has access to a dataset $S=L\cup U$, such that $L=\left\{(x_{i},y_{i})\right\}_{i=1}^{m}$ and $U=\left\{x_{j}\right\}_{j=1}^{u}$.
-In this setting, semi-supervised (SSL) methods attempt to first learn an initial model $f_{init}$ using the labeled data $L$ via a supervised learning algorithm, and in sequence, an attempt is being made in order to harness the information that is hidden in the unlabeled set $U$, so that a better model $f$ can be learnt.
-3.2 Meta Pseudo Labels
-In order to discuss Meta Pseudo Labels, first we need to discuss Pseudo Labels.
-Pseudo Labels (PL) [48] is a broad paradigm for SSL, which involves a teacher and a student in the learning process.
-The teacher looking at the unlabeled set $U$, selects a few instances and assigns pseudo-labels to them.
-Subsequently these pseudo-examples are shown to the student as if they were legitimate examples so that the student can learn a better model.
-In this context, the teacher $f_{\theta_{T}}$ is fixed and the student is trying to learn a better model by aligning its predictions to pseudo-labeled batches $(X_{u},f_{\theta_{T}}(X_{u}))$.
-Thus, on a pseudo-labeled batch, the student optimizes its parameters using
-$\theta_{S}^{\text{PL}}\in\operatorname*{arg\,min}_{\theta_{S}}\ell\left(\operatorname*{arg\,max}_{\xi}f_{\theta_{T}}(X_{u})|_{\xi},f_{\theta_{S}}(X_{u})\right)\,.$ (1)
-The hope is that updated student parameters $\theta_{S}^{\text{PL}}$ will behave well on the labeled data $L=(X_{L},Y_{L})=\left\{(x_{i},y_{i})\right\}_{i=1}^{m}$; i.e., the loss $\mathcal{L}_{L}\left(\theta_{S}^{\text{PL}}\left(\theta_{T}\right)\right)=\ell\left(Y_{L},f_{\theta_{S}^{\text{PL}}}(X_{L})\right)$ is small, where we defined $\mathcal{L}\left(\theta_{S}^{\text{PL}}\left(\theta_{T}\right)\right)$ so that the dependence between $\theta_{S}$ and $\theta_{T}$ is explicit.
-Meta Pseudo Labels (MPL) [59] is exploiting this dependence between $\theta_{S}$ and $\theta_{T}$.
-In other words, because the loss that the student suffers on $\mathcal{L}_{L}\left(\theta_{S}^{\text{PL}}\left(\theta_{T}\right)\right)$ is a function of $\theta_{T}$, and making a similar notational convention for the unlabeled batch $\mathcal{L}_{u}\left(\theta_{T},\theta_{S}\right)=\ell\left(\operatorname*{arg\,max}_{\xi}f_{\theta_{T}}(X_{u})|_{\xi},f_{\theta_{S}}(X_{u})\right)$, then one can further optimize $\theta_{T}$ as a function of the performance of $\theta_{S}$:
-$\left\{\begin{array}[]{rl}\displaystyle\min_{\theta_{T}}&\displaystyle\mathcal{L}_{L}\left(\theta_{S}^{\text{PL}}\left(\theta_{T}\right)\right)\,,\\[3.00003pt] \text{where}&\displaystyle\theta_{S}^{\text{PL}}(\theta_{T})\in\operatorname*{arg\,min}_{\theta_{S}}\mathcal{L}_{u}\left(\theta_{T},\theta_{S}\right)\,.\end{array}\right.$ (2)
-However, because the dependency between $\theta_{T}$ and $\theta_{S}$ is complicated, a practical approximation [51, 35] is obtained via $\theta_{S}^{\text{PL}}\approx\theta_{S}-\eta_{S}\cdot\nabla_{\theta_{S}}\mathcal{L}_{u}\left(\theta_{T},\theta_{S}\right)$ which leads to the practical teacher objective:
-$\min_{\theta_{T}}\mathcal{L}_{L}\left(\theta_{S}-\eta_{S}\cdot\nabla_{\theta_{S}}\mathcal{L}_{u}\left(\theta_{T},\theta_{S}\right)\right)\,.$ (3)
-For more information the interested reader can refer to [59].
-3.3 Co-Training
-Co-training relies on the existence of two different views of the instance space $\mathcal{X}$; that is, $\mathcal{X}=\mathcal{X}^{(1)}\times\mathcal{X}^{(2)}$.
-Hence, an instance $x\in\mathcal{X}$ has the form $x=(x^{(1)},x^{(2)})\in\mathcal{X}^{(1)}\times\mathcal{X}^{(2)}$.
-We say that the views $x^{(1)}$ and $x^{(2)}$ are complementary.
-Co-training requires two assumptions in order to work.
-First, it is assumed that, provided enough many examples, each separate view is enough for learning a meaningful model.
-The second is the conditional independence assumption [12], which states that $x^{(1)}$ and $x^{(2)}$ are conditionally independent given the label of the instance $x=(x^{(1)},x^{(2)})$.
-Thus, with co-training, two models $f^{(1)}$ and $f^{(2)}$ are learnt, which help each other using different information that they capture from the different views of each instance.
-More information is presented below.
-Given a dataset $S=L\cup U$ (Section 3.1 has notation), we can have access to two different views of the dataset due to the natural partitioning of the instance space $\mathcal{X}$; that is, $S^{(1)}=L^{(1)}\cup U^{(1)}$ and $S^{(2)}=L^{(2)}\cup U^{(2)}$.
-Initially two models $f^{(1)}$ and $f^{(2)}$ are learnt using a supervised learning algorithm based on the labeled sets $L^{(1)}$ and $L^{(2)}$ respectively.
-Learning proceeds in iterations, where in each iteration a number of unlabeled instances are integrated with the labeled set by assigning pseudo-labels to them that are predicted by $f^{(1)}$ and $f^{(2)}$, so that improved models $f^{(1)}$ and $f^{(2)}$ can be learnt by retraining on the augmented labeled sets.
-In particular, in each iteration, each model $f^{(v)}$ predicts a label for each view $x^{(v)}\in U^{(v)}$.
-The top-$k$ most confident predictions by model $f^{(v)}$ are used to provide a pseudo label for the respective complementary instances and then these pseudo-labeled examples are integrated in to the labeled set for the complementary view.
-In the next iteration a new model is trained based on the augmented labeled set.
-Both views of the used instances are dropped from $U$ even if only one of the views may be used for augmenting the appropriate labeled set.
-This process repeats until the unlabeled sets $U^{(v)}$ are exhausted, yielding models $f^{(1)}$ and $f^{(2)}$ corresponding to the two views.
-Co-Training is evaluated using the joint prediction of these models which is the re-normalized element-wise product of the two predictions:
-$\frac{f^{(1)}\left(x^{(1)}\right)\otimes f^{(2)}\left(x^{(2)}\right)}{\left\lVert f^{(1)}\left(x^{(1)}\right)\otimes f^{(2)}\left(x^{(2)}\right)\right\rVert_{1}}\,.$ (4)
-Co-training was introduced by Blum and Mitchell [12] and their paradigmatic example was classifying web pages using two different views: (i) the bags of words based on the content that they had and (ii) the bag of words formed by the words that were used for hyperlinks pointing to these webpages.
-In our context the two different views are obtained via different embeddings that we obtain when we use images as inputs to different foundation models and observe the resulting embeddings.
-4 Proposed Methods
-4.1 View Construction
-Past methods of view construction include manual feature subsetting [33], automatic feature subsetting [22, 65], random feature subsetting [15, 16], random subspace selection [16], and adversarial examples [60].
-We propose a new method of view construction which is motivated by modern approaches to semi-supervised and representation learning in computer vision and natural language processing: pretext learning.
-Given two pretext tasks we can train a model to accomplish each task, then use their learned embedding spaces as separate views for classification.
-If we have chosen our pretext tasks well, then the representations will be different, compressed, and capture sufficient information to perform the desired downstream task.
-While our approach is more widely applicable, it is particularly well-suited to CV.
-There are a plethora [38, 23, 24, 19, 49, 20, 61, 58, 42] of competitive learned representations for images.
-Recently it has been popular to treat these representations as task-agnostic foundation models.
-We show that the existence of these models eliminates the need to train new models for pretext tasks to achieve state-of-the-art performance for SSL-classification.
-Furthermore, it is comparatively very cheap to use the foundation models instead as not only can we forego training on a pretext task, but we are able to completely bypass the need to train a large convolutional neural network (CNN) or vision transformer (ViT).
-In fact, we compute the embedding for each of the images in our dataset only one time (as opposed to applying some form of data augmentation during training and making use of the embedding process during training).
-So, at training time the only weights we need to use at all are those in the model we train on top of the learned representation.
-This makes our approach incredibly lightweight compared to traditional CNN/ViT training procedures.
-4.2 Meta Co-Training
-In addition to our method of view construction we propose a novel algorithm for co-training which is based on Meta Pseudo Labels [59], called Meta Co-Training (MCT).
-Algorithm 1 presents the details of this method.
+That is, the learning algorithm has access to a dataset $S=L\cup U$, such that $L=\left\{(x_{i},y_{i})\right\}_{i=1}^{m}$ and $U=\left\{x_{j}\right\}_{j=1}^{u}$.
+For $L$ the instance part is denoted as $X_{L}$ and the label part is denoted as $Y_{L}$.
+In this setting, semi-supervised learning (SSL) methods attempt to first learn an initial model $f_{init}$ using the labeled data $L$ via a supervised learning algorithm, and in sequence, an attempt is being made in order to harness the information that is hidden in the unlabeled set $U$, so that a better model $f$ can be learnt.
+View Construction.
+A view, in the context of co-training methods, is a subset of the input features.
+A view typically has two properties: (i) it is sufficient for predicting the desired quantity, and (ii) it is conditionally independent of other views given the label.
+Previous methods of view construction for co-training include manual feature subsetting [19], automatic feature subsetting [11], random feature subsetting [7], random subspace selection [8], and adversarial examples [44].
+We use self-supervised learning to construct views for our experiments.
+However, if one has access to a dataset where the instances have two different views, then one can still obtain useful representations with this approach using the same self-supervised model or none at all.
+While our approach is more widely applicable, it is particularly well-suited to computer vision.
+There are a plethora [23, 12, 10, 34, 45, 42, 26] of competitive learned representations for images.
+Choosing foundation model representations as views makes our approach lightweight enough to be feasible with limited hardware resources.
+3 Proposed Method: Meta Co-Training
+We propose a novel semi-supervised learning algorithm called Meta Co-Training (MCT), which is described below.
+MCT operates within a bi-level student-teacher optimization framework.
+A student is optimized to replicate the labels its teacher gives to a set of unlabeled points while a teacher is optimized to provide labels to the same unlabeled points such that its student learns to predict well on labeled data.
+We call MCT a co-training algorithm because it operates on two views of the data, however unlike the original co-training method [5] we do not train multiple models for each view sequentially.
+Each view has exactly one model which acts as both a teacher for the other view and a student for its own view.
+That is to say that the model learning on view 1 will take an unlabeled instance from view 1 as input and provide the pseudo-label prediction to the student for the same unlabeled instance in view 2.
+The student – the model learning on view 2 – will use the pseudo-label to learn to predict similarly.
+The lower of the two levels of the optimization is the student optimization (Equation 2). The student parameters $\theta_{S}$ are optimized as a function of the teacher parameters $\theta_{T}$:
+$\mathcal{L}_{u}\left(\theta_{T},\theta_{S}\right)=\ell\left(\operatorname*{arg\,max}_{\xi}f_{\theta_{T}}(U)|_{\xi},f_{\theta_{S}}(U)\right)$ (1)
+$\theta_{S}^{\prime}=\theta_{S}^{\text{PL}}(\theta_{T})\in\operatorname*{arg\,min}_{\theta_{S}}\mathcal{L}_{u}\left(\theta_{T},\theta_{S}\right)\,.$ (2)
+The teacher optimization (Equation 4) forms the upper level of the optimization:
+$\mathcal{L}_{L}\left(\theta_{S}^{\prime}\right)=\ell\left(Y_{L},f_{\theta_{S}^{\prime}}(X_{L})\right)$ (3)
+$\theta_{T}^{\prime}=\operatorname*{arg\,min}_{\theta_{T}}\displaystyle\mathcal{L}_{L}\left(\theta_{S}^{\prime}\right)\,.$ (4)
+All together the objective of MCT from the perspective of the current view model as the teacher is
+$\min_{\theta_{T}}\mathcal{L}_{u}\left(\theta_{T},\theta_{S}\right)+\mathcal{L}_{L}\left(\theta_{S}^{\prime}\right)\,.$ (5)
+It is cumbersome to speak of teachers and students when each model is both.
+We give the objective of MCT in Equation 6 as jointly optimizing the following objectives with $f_{\theta_{1}}$ the model learning on view 1 and $f_{\theta_{2}}$ the model learning on view 2.
+$\min_{\theta_{1},\theta_{2}}\mathcal{L}_{u}\left(\theta_{1},\theta_{2}\right)+\mathcal{L}_{L}\left(\theta_{2}^{\prime}\right)\ +\mathcal{L}_{u}\left(\theta_{2},\theta_{1}\right)+\mathcal{L}_{L}\left(\theta_{1}^{\prime}\right)\,$ (6)
+Co-training algorithms are traditionally evaluated using the joint prediction of the constituent models; i.e., the re-normalized element-wise product of the individual predictions.
+Figure 1: At each step $t\in\{1,\ldots,T\}$ of meta co-training the models that correspond to the so-far learnt parameters $\theta_{1;t}$ and $\theta_{2;t}$ play the role of the student and the teacher simultaneously using batches for their respective views.
+Pseudo-labeling occurs on complementary views so that the teacher can provide the student with labels on an unlabeled batch.
+Labeled batches may, or may not, use complementary views as the purpose that they serve is to calculate the risk of the student model on the labeled batch and this result signals the teacher model to update its weights accordingly.
+To make tractable the optimization in Equation 6 we approximate the student and teacher optimizations using alternating iterations of stochastic gradient descent.
+In Algorithm 1 the student updates are in lines 13-14 while the teacher updates are in lines 24-25.
+An overview of MCT is provided in Figure 1.
+At each step $t\in\{1,\ldots,T\}$ of meta co-training the models that correspond to the so-far learnt parameters $\theta_{1;t}$ and $\theta_{2;t}$ participate in a student-teacher framework in which each model plays the role of the student for their view and the teacher for the other model simultaneously.
+The representations obtained from the view construction process of Section 2 form different views that are used by MCT.
+By training two models on two views to collaborate we can take advantage of the aspects of MPL and Co-Training that make each algorithm successful.
+Co-training is more effective when the two constituent models fail (predict incorrectly) independently.
+During MCT the teacher model will receive positive feedback precisely when it makes a prediction that the other model would not have made which results in an improvement.
+MCT trains models in a way which encourages each teacher to express independence while it teaches each student to incorporate the improvement.
+Naturally, once the two models predict similarly then the gradient of the student parameters with respect to the teacher’s labels is zero and they both cease to improve.
+This, however, is also when MPL ceases to improve and it is unlikely to have models which are as diverse as those which are trained from independent views.
+Algorithm 1 Meta Co-Training (Section 2 has notation)
 1:  Input: $L^{(1)},L^{(2)},U^{(1)},U^{(2)},f^{(1)},f^{(2)},\theta_{1},\theta_{2},T,\ell,\eta$
@@
 5:     $U_{t}^{(2)}\leftarrow getOtherView(U_{t}^{(1)},2)$
-6:     // Predict soft pseudo labels
+6:     // Predict soft pseudo-labels
 7:     $\hat{y}^{(1)}\leftarrow f^{(1)}_{\theta_{1}}(U_{t}^{(1)})$
 8:     $\hat{y}^{(2)}\leftarrow f^{(2)}_{\theta_{2}}(U_{t}^{(2)})$
-9:     // Sample the pseudo labels from the discrete distribution over the classes
+9:     // Sample the pseudo-labels from the discrete distribution over the classes
 10:     $PL^{(1)}\sim\hat{y}^{(1)}$
 11:     $PL^{(2)}\sim\hat{y}^{(2)}$
-12:     // MPL student update applied to both models
+12:     // MCT student update applied to both models
 13:     $\theta^{\prime}_{1}\leftarrow\theta_{1}-\eta\cdot\nabla_{\theta_{1}}\ell(PL^{(2)},\hat{y}^{(1)})$
@@
 17:     $X_{t}^{(2)},Y_{t}^{(2)}\leftarrow SampleBatch(L^{(2)})$
-18:     // MPL teacher update applied to both models
+18:     // MCT teacher update applied to both models
 19:     $\hat{y}^{\prime(1)}\leftarrow f^{(1)}_{\theta^{\prime}_{1}}(X_{t}^{(1)})$
@@
 26:  end for
-Algorithm 1 Meta Co-Training (Section 3.1 has notation)
-At each step $t$ of the algorithm each model is first updated based on the pseudo labels the other has provided on a batch of unlabeled data (Lines 16-17).
-Then each model is updated to provide labels that encourage the other to predict more correctly on the labeled set based on the performance of the other (Lines 23-24).
-We refer the reader to Section 3.2 for a brief explanation as to why these updates make sense, or to [59] for a more complete discussion of the MPL approach.
-Geometrically, we can understand this as updating the parameters in the direction of increasing confidence on our pseudo labels if those labels helped the other model perform better on the labeled set, or in the opposite direction if it hurt the other model’s performance on the labeled set.
-Note that contrary to traditional co-training, no new pseudo-labeled data are added to the labeled set, neither are any data removed from the unlabeled set after each step.
-5 Experiments
-Towards evaluating the proposed methods of view construction and co-training we first constructed views of the data utilizing five different learned representations.
-Our primary comparisons were on the widely-used ImageNet [31] classification task; we provide experimental results on additional image classification tasks.
-To produce the views used to train the classifiers during co-training and meta co-training we used the embedding space of five representation learning architectures: The Masked Autoencoder (MAE) [42], DINOv2 [58], SwAV [19], EsViT [49], and CLIP [61].
-We selected the models which produce the views as they have been learned in an unsupervised way, have been made available by the authors of their respective papers for use in PyTorch, and have been shown to produce representations that are appropriate for ImageNet classification.
+Algorithm Description.
+Algorithm 1 presents pseudocode for the algorithm.
+The function SampleBatch, is sampling a batch from the dataset.
+The function getOtherView returns the complementary view of the first argument; the second argument clarifies which view should be returned (“2” in line 5).
+At each step $t$ of the algorithm each model is first updated based on the pseudo-labels the other has provided on a batch of unlabeled data (lines 13-14).
+Then each model is updated to provide labels that encourage the other to predict more correctly on the labeled set based on the performance of the other (lines 24-25).
+Further discussion of these update rules can be found in Appendix A.
+Each student model evaluates the performance of their new weights on separate labeled batches.
+This performance provides feedback to the teacher model.
+Geometrically, we can understand this as updating the parameters in the direction of increasing the teacher’s confidence on pseudo-labels, assuming those pseudo-labels helped the student model perform better on the labeled set.
+Along these lines, the teacher parameters are updated in the opposite direction if it hurt the student’s performance on the labeled set.
+In general MCT does not require any preexisting method of view construction to be applied.
+If a dataset naturally provides two views, then there is no need to construct additional views.
+One can use their favorite pre-trained representation or none at all and then apply MCT as described above.
+Additionally we can relax the assumption that there is correspondence between the labeled examples for each view.
+If it happens to be the case that it is easy to collect unlabeled examples from each view which correspond to the same label but existing labeled sets for each view do not, then MCT can still be applied.
+This is because the labeled data used to evaluate the change in the performance of the student does not need to have corresponding instances in the teacher’s view.
+In all of our experiments in this work we use preexisting foundation models to construct views for our co-training methods, and thus in all cases there exist two views for each labeled example.
+In Algorithm 1 it is illustrative to consider all versions of the parameters within a step and all of their gradients to exist simultaneously in memory.
+This is undesirable in practice given the large size of common neural networks and their gradients.
+Instead, we compute $\nabla_{\theta_{2}}\ell(PL^{(2)},\hat{y}^{(2)})$ and $\nabla_{\theta_{1}}\ell(PL^{(1)},\hat{y}^{(1)})$ (used in lines 24-25) first, then we compute (and apply) $\nabla_{\theta_{1}}\ell(PL^{(2)},\hat{y}^{(1)})$ and $\nabla_{\theta_{2}}\ell(PL^{(1)},\hat{y}^{(2)})$ (used first in lines 13-14 and reused in lines 21-22), then we compute $\nabla_{\theta^{\prime}_{1}}\ell(Y_{t}^{(1)},\hat{y}^{\prime(1)})^{T}$ and $\nabla_{\theta^{\prime}_{2}}\ell(Y_{t}^{(2)},\hat{y}^{\prime(2)})^{T}$ which we subsequently use to compute $h^{(1)}$ and $h^{(2)}$.
+We finish the iteration with lines 24-25 using the saved gradient from the first step.
+With this approach we only have to compute the forward pass and gradient for the student update once, and we only keep one version of the model weights in memory at a time.
+Relation to Ensemble Methods.
+As our approach utilizes multiple models which collaborate during prediction, we acknowledge that it bears similarity to the family of ensemble methods.
+Typically, co-trained models are evaluated against other SSL approaches.
+In our evaluation we compare our semi-supervised method to supervised deep ensembles to illustrate the impact of MCT versus ensemble methods.
+We show that when MCT works well, the performance benefit cannot be attributed to simply utilizing multiple models.
+4 Experimental Evaluation
+Towards evaluating our proposed method we compare to other self-supervised and semi-supervised learning algorithms using the ImageNet [17] classification benchmark, as well as Flowers102 [41], Food101 [6], FGVCAircraft [39], iNaturalist [29], and iNaturalist 2021 [28] datasets.
+To produce the views used to train the classifiers during CT and MCT we used the embedding space of five representation learning architectures: The Masked Autoencoder (MAE) [26], DINOv2 [42], SwAV [10], EsViT [34], and CLIP [45].
+We selected the models which produce the views as they have been learned in an unsupervised way, have been made available by the authors of their respective papers for use in PyTorch, and have been shown to produce representations that are appropriate for computer vision classification tasks.
+Hyperparameters used during training are included in Table 11 in the appendix for CT, MCT, and deep ensembles.
+Table 1 provides a description of the different datasets used for training and testing, as well as the number of classes for each dataset.
+Note that in the 10% experiments on the Flowers102 dataset each class has only 1 label.
+All datasets are approximately class-balanced with the exception of the iNaturalist dataset.
+In all cases we maintain the original class distribution when sampling subsets.
+For ImageNet we use the split published with the SimCLRv2 [13] repository.
+All subsets are created with seeded randomness of a common seed (13) which ensures a fair comparison between methods.
+Table 1: Characteristics of datasets used.
+Dataset #train #test #classes
+FGVCAircraft $3$ $3$ 100
+Flowers102 1020 6149 102
+Food101 75,750 25,250 101
+iNaturalist 175,489 29,083 1010
+ImageNet 1,281,167 50,000 1000
+4.1 Experimental Evaluation on ImageNet
+As we discussed in Section 2, CT relies on two assumptions about the sufficiency and independence of the two views in our data.
+We conducted experiments to verify these assumptions.
+Below we give a summary of our results.
 On the Sufficiency of the Views.
-Motivated by the analysis and experiments of [34] we made an effort to verify the sufficiency and independence of the representations learned by the aforementioned models.
 Sufficiency is fairly easy to verify.
-If we choose a reasonable sufficiency threshold for the task, for ImageNet say $75\%$ top-1 accuracy, then we can train simple models on the views of the dataset we have constructed and provide examples of functions which demonstrate the satisfaction of the property.
-We tested a single linear layer with softmax output (Table 1), and a simple 3-layer multi-layer perceptron (MLP) with 1024 neurons per layer and dense skip connections (previous layer outputs were concatenated to input for subsequent layers) (Table 2) to provide a lower bound on the sufficiency of the views on the standard subsets of the ImageNet labels for semi-supervised classification.
+If we choose a reasonable sufficiency threshold for the task, for ImageNet say close or above $75\%$ top-1 accuracy, then we can train simple models on the views of the dataset we have constructed and provide examples of functions which demonstrate the satisfaction of the property.
+We tested a single linear layer with softmax output to provide a lower bound on the sufficiency of the views on the standard subsets of the ImageNet labels for semi-supervised classification.
+A simple 3-layer multi-layer perceptron (MLP) with 1024 neurons per layer was evaluated (Table 2).
+These were the same models as those used for CT and MCT.
+From these experiments, whose results are shown in Table 2 we can see that if we were to pick a threshold for top-1 accuracy around $75\%$, then CLIP and DINOv2 provide sufficient views for both the $1\%$ and $10\%$ subsets.
+Table 2: MLP performance on individual views. Top-1 accuracy on different subsets of the ImageNet data are shown.
+Model 1% 10%
+MAE 23.4 48.5
+DINOv2 78.4 82.7
+SwAV 12.5 32.6
+EsViT 71.3 75.8
+CLIP 75.2 80.9
+On the Independence of the Views.
+To test the independence of the views generated by the representation learning models, we trained an identical MLP architecture to predict each view from each other view; Table 3 presents our findings, explained below.
+Table 3: Pairwise translation performance of linear probe on the ImageNet dataset.
+A linear classifier is trained on the output of an MLP which is trained to predict one view (columns) from another view (rows) by minimizing MSE.
+The top-1 accuracy (%) of the linear classifier is reported.
+Both the MLP and the linear classifier have access to the entire embedded ImageNet training set.
+MAE DINOv2 SwAV EsViT CLIP
+MAE - 0.139 0.142 0.137 0.129
+DINOv2 0.110 - 0.132 0.135 0.130
+SwAV 0.116 0.147 - 0.137 0.126
+EsViT 0.112 0.140 0.116 - 0.135
+CLIP 0.110 0.146 0.136 0.156 -
+Given as input the view to predict from (rows of Table 3) the MLPs were trained to reduce the mean squared error (MSE) between their output and the view to be predicted (columns of Table 3).
+We then trained a linear classifier on the outputs generated by the MLP given every training set embedding from each view.
+Had the model faithfully reconstructed the output view given only the input view then we would expect the linear classifier to perform similarly to the last column of Table 4.
+The linear classifiers never did much better than a random guess on the ImageNet class achieving at most 0.156% accuracy.
+Thus, while we cannot immediately conclude that the views are independent, it is clearly not trivial to predict one view given any other.
+We believe that this is compelling evidence for the independence of different representations.
+Table 4: Linear probe evaluation for views. Top-1 accuracy on different subsets of the ImageNet data are shown.
 Model 1% 10% 100%
@@
 CLIP 74.1 80.9 85.4
-Table 1: Linear probe evaluation for views. Top-1 accuracy on different subsets of the ImageNet data are shown.
-Model 1% 10%
-MAE 23.4 48.5
-DINOv2 78.4 82.7
-SwAV 12.5 32.6
-EsViT 71.3 75.8
-CLIP 75.2 80.9
-Table 2: MLP performance on individual views. Top-1 accuracy on different subsets of the ImageNet data are shown.
-From these experiments we can see that if we were to pick the $75\%$ top-1 accuracy threshold, which we believe indicates considerable skill without presenting too high of a bar, then CLIP and DINOv2 provide sufficient views for both the $1\%$ and $10\%$ subsets.
-On the Independence of the Views.
-To test the independence of the views generated by the representation learning models, we trained an identical MLP architecture to predict each view from each other view; Table 3 presents our findings, explained below.
-MAE DINOv2 SwAV EsViT CLIP
-MAE - 0.139 0.142 0.137 0.129
-DINOv2 0.110 - 0.132 0.135 0.130
-SwAV 0.116 0.147 - 0.137 0.126
-EsViT 0.112 0.140 0.116 - 0.135
-CLIP 0.110 0.146 0.136 0.156 -
-Table 3: Pairwise translation performance of linear probe on the ImageNet dataset.
-A linear classifier is trained on the output of an MLP which is trained to predict one view (columns) from another view (rows) by minimizing MSE.
-The top-1 accuracy (%) of the linear classifier is reported.
-Both the MLP and the linear classifier have access to the entire embedded ImageNet training set.
-Given as input the view to predict from (rows of Table 3) the MLPs were trained to reduce the mean squared error (MSE) between their output and the view to be predicted (columns of Table 3).
-We then trained a linear classifier on the outputs generated by the MLP given every training set embedding from each view.
-Had the model faithfully reconstructed the output view given only the input view then we would expect the linear classifier to perform similarly to the last column of Table 1.
-The linear classifiers never did much better than a random guess on the ImageNet class achieving at most 0.156% accuracy.
-Thus, while we cannot immediately conclude that the views are independent, it is clearly not trivial to predict one view given any other.
-We believe that this is compelling evidence for the independence of different representations.
-Co-Training Experiments on ImageNet.
-Having constructed at least two strong views of the data, and suspecting that these views are independent we hypothesize that co-training will work on this data.
-In Table 3(a) we show the results of applying the standard co-training algorithm [12].
-At each iteration of co-training the model made a prediction for all of the instances in $U$, following (4).
-We assigned labels to the 10% of the total original unlabeled data with most confident predictions for both models.
-When confident predictions conflicted on examples, we returned those instances to the unlabeled set, otherwise they entered the labeled set of the other model with the assigned pseudo-label.
-Comparing co-training accuracy on ImageNet 10% (resp.
-1%) shown in Table 3(a) compared to MLP accuracy shown in Table 2, we observe that co-training top-1 accuracy is on average 18.7% (resp.
-16.8%) higher than MLP accuracy.
-Moreover, co-training top-1 accuracy is 2.4% (resp.
-1.7%) higher overall.
-In other words, co-training resulted in a significant increase in prediction accuracy but not after any iterations of co-training have occurred; see Figures 1 and 2.
-Furthermore, predicting according to (4) yielded better accuracy than the predictions of each individual model.
-In addition, in our experiments on ImageNet, as co-training iterations proceeded top-1 accuracy decreased.
-Later we show that this was not always the case (see Figure 9 in the appendix).
-The decrease was less pronounced when the views perform at a similar level (see Figure 3).
-Despite the poor performance of the pseudo-labeling during co-training, just computing the re-normalized product of the two best performing views performed as well as the previous SotA method for semi-supervised classification on the ImageNet-10% task (see Co-Training in Table 5).
-The performance of co-training as a method was rather disappointing in that the algorithm is unable to leverage pseudo-labels to improve the accuracy on this task; however, the results do show that our view-construction method has merit.
-Given the difficulty of translating between the views, and the fact that the accuracy yielded by (4) is greater than its constituent views, we believe the poor performance of the pseudo-labeling step in co-training was due to the insufficiency of at least one view rather than view-dependence.
-Figure 1: Top-1 accuracy of co-training iterations for the ImageNet 10% dataset. The average over all 10 pairs, the best pair, and the worst pair at each iteration are shown.
-Figure 2: Top-1 accuracy of co-training iterations for the ImageNet 1% dataset. The average over all 10 pairs, the best pair, and the worst pair at each iteration are shown.
-Figure 3: Top-1 accuracy of co-training iterations on the CLIP and DINOv2 views for the ImageNet 10% dataset.
-Figure 4: Top-1 accuracy of co-training iterations on the CLIP and DINOv2 views for the ImageNet 1% dataset.
-DINOv2 SwAV EsViT CLIP
-MAE 81.8 (75.5) 30.5 (11.4) 77.1 (66.2) 78.8 (66.8)
-DINOv2 78.5 (74.5) 83.3 (78.9) 85.1 (80.1)
-SwAV 70.6 ( 64.9) 75.5 (67.4)
-EsViT 82.3 ( 76.9)
-(a) Co-Training
-DINOv2 SwAV EsViT CLIP
-MAE 81.2 (73.8) 37.6 (8.1) 73.6 ( 63.8) 78.6 (63.5)
-DINOv2 77.3 ( 70.7) 83.4 (79.2) 85.2 (80.7)
-SwAV 74.4 (67.3) 77.1 (67.4)
-EsViT 82.4 (77.5)
-(b) Meta Co-Training
-Table 4: Co-training and meta co-training top-1 accuracy of view combinations on 10% (1%) of ImageNet labels.
-As co-training does not depend on the order of the views, we show only upper-diagonal entries of the pairwise comparison.
-Diagonal entries would correspond to views which are trivially dependent.
-Note that all $\binom{5}{2}=10$ different combinations of views are shown in each table.
+Having constructed at least two strong views of the data, and suspecting that these views are independent we hypothesize that MCT and CT will work on this data.
+Co-Training Baseline Experiments.
+At each iteration of CT the models made predictions for all instances in $U$.
+We assigned labels to the 10% of the most confident predictions in $U$ for both models.
+When confident predictions conflicted on examples, we returned those instances to the unlabeled set, otherwise their complementary views were entered into the labeled set of the other model with the assigned pseudo-label.
+For CT, joint predictions yielded better accuracy than the predictions of each individual model.
+In our experiments on ImageNet, as CT iterations proceeded, top-1 accuracy decreased.
+Later we show that this was not always the case (see Figure 4).
+The decrease was less pronounced when the views perform at a similar level; see Figure 2.
+In contrast MCT does not exhibit performance degradation after warmup.
+Figures 2 and 3 show results on ImageNet 10%.
+We see that while the CT performance decreases after the supervised warmup, MCT performance increases.
+Despite the poor performance of the pseudo-labeling during CT, using the two best performing views CT performed as well as the previous SoTA method for SSL classification on the ImageNet-10%; see Co-Training in Table 5.
+CT as a method was rather disappointing in that the algorithm is unable to leverage pseudo-labels to improve the accuracy on this task.
+Given the difficulty of translating between the views, and the fact that the accuracy yielded by the joint prediction is greater than its constituent views, we believe the poor performance of the pseudo-labeling step in CT was due to imbalanced information content between views rather than view-dependence.
+Figure 2: Top-1 accuracy of CT iterations on the CLIP and DINOv2 views for the ImageNet 10% dataset.
+Figure 3: MCT using the CLIP and DINOv2 views as a function of the training step. Models are trained on 10% of the ImageNet labels.
+Table 5: Performance of different approaches on ImageNet dataset.
+An asterisk (*) indicates models that were trained on top of embeddings generated by much larger models.
+During training we do not need to alter the parameters of these larger models, and these larger models need not see data used for downstream classification during their training, however training of downstream models would not have been possible without the embeddings.
+Models on the lower half of the table use unlabeled data during classification training.
 Reference Model Method ImageNet-1% ImageNet-10%
@@
 [44] Deep Co-Training (ResNet-18) Co-Training - 53.5
+[51] UDA (ResNet50) UDA - 68.78
+[49] FixMatch (ResNet-50) FixMatch - 71.46
+[43] MPL (EfficientNet-B6-Wide) MPL - 73.9
 [9] Semi-ViT (ViT-L) Self-trained 77.3 83.3
 [37] REACT (ViT-L) REACT 81.6 85.1
-[59] MPL (EfficientNet-B6-Wide) MPL - 73.9
 [5] Co-Training (MLP)* Co-Training 80.1 85.1
-MCT (MLP)* - ours MCT 80.7 85.8
-Table 5: Performance of different approaches on ImageNet dataset.
-An asterisk (*) indicates models that were trained on top of embeddings generated by much larger models.
-During training we do not need to alter the parameters of these larger models, and these larger models need not see data used for downstream classification during their training, however training of downstream models would not have been possible without the embeddings.
-Models on the lower half of the table use unlabeled data during classification training.
-Meta Co-Training Experiments on ImageNet.
-To draw fair comparisons we fix the model architecture and view set from our experiments on co-training.
-As in [59] a supervised loss is optimized jointly with the MPL loss.
-We found that beginning with a warmup period in which the model trained in a strictly supervised way expedited training.
-Unlike MPL, we did not use Unsupervised Data Augmentation [75], because we embedded the dataset only once, and due to the nature of the pretext tasks that were used to generate the views we did not believe data augmentation would have much of an effect.
-Details of our training recipe appear in Table 10.
-Comparing meta co-training accuracy on ImageNet 10% (resp.
-1%) shown in Table 3(b) compared to MLP accuracy shown in Table 3(a), we observe that meta co-training top-1 accuracy yielded a difference on average 0.7% (resp.
--0.6%) compared to co-training top-1 accuracy.
-Moreover, meta co-training top-1 accuracy is 0.1% (resp.
-0.6%) higher overall.
-Note that on all view pairs in which co-training beats meta co-training, with the exception of the $10\%$ (EsViT, MAE) pair, one of the views alone was better than co-training.
-The single DINOv2 view outperforms all such view pairs.
-These are all cases in which one of the views was significantly weaker than the other.
-For the experiments in which we had two strong (e.g., not SwAV or MAE) views, meta co-training always had better performance than co-training.
+[32] Deep Ensemble * Deep Ensemble 80.0 84.3
+Meta Co-Training (MLP)* - ours Meta Co-Training 80.7 85.8
+Meta Co-Training Experiments.
+To draw fair comparisons, we fix the model architecture and view set from our experiments on CT.
+As in MPL [43] a supervised loss is optimized jointly with a loss on pseudo-labels.
+We found that beginning with a warmup period in which the models for each view are trained in a strictly supervised way expedited training.
+This warmup period occurs for the same number of updates as the first training phase of CT, so each method starts with the same number of supervised updates before pseudo-labeling.
+Unlike MPL, we did not use Unsupervised Data Augmentation [51], because we embedded the dataset only once, and using self-supervised models to generate the views we did not believe data augmentation would have much of an effect.
+Details of our training recipe are given in Table 11.
+Table 6 shows the results of MCT on all of the possible combinations of views we used.
+With one exception, the better the performance of the views, the better MCT performs compared to CT.
+In nearly all cases in which both views are strong, MCT performs better than CT.
+These are also the cases with the highest performance overall.
+Comparing CT accuracy on ImageNet 10% (resp.
+1%) shown in Table 6(a) compared to MLP accuracy shown in Table 2, we observe that CT top-1 accuracy is on average 18.7% (resp.
+16.8%) higher than MLP accuracy.
+Moreover, the best CT top-1 accuracy is 2.4% (resp.
+1.7%) higher compared to the best top-1 MLP accuracy.
+In Table 6(c) we show the performance of an ensemble of models trained on concatenated view pairs but we provide more information on the comparison between MCT and deep ensembles in a separate paragraph below.
+Table 6: CT, MCT, and ensemble top-1 accuracy of view combinations on 10% (1%) of ImageNet labels.
+As CT and MCT do not depend on the order of the views, we show only upper-diagonal entries of the pairwise comparison.
+All $\binom{5}{2}=10$ different combinations of views are shown in each table.
+In the case of the MLP ensemble the order of the pairs of views also does not matter.
+These views are concatenated in order to form a larger unified view and hence different order of the views would correspond to different order of the attributes in the embeddings, which does not change their information content.
+(a) Co-Training
+DINOv2 SwAV EsViT CLIP
+MAE 81.8 (75.5) 30.5 (11.4) 77.1 (66.2) 78.8 (66.8)
+DINOv2 78.5 (74.5) 83.3 (78.9) 85.1 (80.1)
+SwAV 70.6 ( 64.9) 75.5 (67.4)
+EsViT 82.3 ( 76.9)
+(b) Meta Co-Training
+DINOv2 SwAV EsViT CLIP
+MAE 81.2 (73.8) 37.6 (8.1) 73.6 ( 63.8) 78.6 (63.5)
+DINOv2 77.3 ( 70.7) 83.4 (79.2) 85.2 (80.7)
+SwAV 74.4 (67.3) 77.1 (67.4)
+EsViT 82.4 (77.5)
+(c) MLP Ensemble
+DINOv2 SwAV EsViT CLIP
+MAE 83.0 (78.9) 38.1 (17.9) 75.0 ( 72.1) 79.4 (76.1)
+DINOv2 83.7 ( 79.1) 81.2 (78.5) 84.2 (80.0)
+SwAV 74.4 (72.5) 81.2 (76.8)
+EsViT 78.7 (73.5)
+MCT and CT are compared on ImageNet 10% (resp., 1%) in Table 5.
+We observe that MCT top-1 accuracy is 0.7% (resp., 0.6%) higher overall.
+An ablation study for the views is given in Table 6.
+On all view pairs in which MCT or CT provide no benefit over a single view are cases in which one of the views was significantly weaker than the other.
+This makes sense as one model has nothing to learn from the other.
+For the experiments in which we had two strong views, MCT always outperforms CT.
+Comparison to Ensemble Methods.
+Ensemble methods are common in practice to boost the performance of supervised models.
+To show that our method provides benefit outside of just adding an additional model, we compare to a Deep Ensemble [32] of five models for each individual view.
+In Table 7 we show the performance of ensembles trained on the individual views.
+We show the performance of an ensemble trained on each view pair concatenated in Table 6(c).
+We show the performance of an ensemble trained on the concatenation of all four views we used to achieve our best performing MCT experiment in Table 8.
+In some cases the CT models perform better than the ensemble of models which all have the same input for each sample (the concatenation of the CT views).
+We believe that CT before pseudo-labeling (which is essentially an ensemble of two different views) out-performs the deep ensembles because the CT models are fewer but they are more diverse.
+The CT models are more diverse as they are trained on different views, while the ensemble models are trained on the concatenation of the views.
+The concatenation has the same information but may not produce meaningfully different models in an ensemble.
+Ensuring diversity of ensemble members is an important problem for ensembling methods, so it seems likely that a less diverse ensemble would perform worse.
+If this is indeed the case, then it would support our hypothesis that CT is failing to improve after pseudo-labeling due to an imbalance in the sufficiency of views and not a lack of independence.
+MCT provides performance gains over CT and the deep ensembles, so we can be confident that these gains cannot be attributed solely to the use of multiple models.
+Table 7: MLP ensemble performance on individual views. Top-1 accuracy on different subsets of the ImageNet data are shown.
+Model 1% 10%
+MAE 1.7 3.8
+DINOv2 79.2 82.8
+SwAV 19.7 40.3
+EsViT 72.1 75.0
+CLIP 76.4 80.8
+4.2 Stronger Views by Concatenation
 Finally, we take the four views which had the greatest performance and constructed two views out of them by concatenating them together.
-We constructed one view CLIP |||| EsViT and the second DINOv2 |||| SwAV and measured the performance of co-training and meta co-training on these views in Table 6.
-Interestingly, co-training does not benefit from these larger views; see Table 6.
+We constructed one view CLIP |||| EsViT and the second DINOv2 |||| SwAV and measured the performance of CT and MCT on these views in Table 8.
+Interestingly, CT does not benefit from these larger views; see Table 8.
+Table 8: Co-training and MCT evaluated on the ImageNet dataset using the CLIP |||| EsViT and DINOv2 |||| SwAV views. The Deep Ensemble is evaluated on the concatenation of all four views.
 Method 1% 10%
+Deep Ensemble 80.0 84.3
 Co-Training 80.0 84.8
-MCT 80.5 85.8
-Table 6: Meta co-training and co-training evaluated on the ImageNet dataset using the CLIP |||| EsViT and DINOv2 |||| SwAV views.
+Meta Co-Training 80.5 85.8
 The additional information may have allowed the individual models to overfit their training data quickly.
-During meta co-training we observed that after one model reaches 100% training accuracy, its validation accuracy can still improve by learning to label for the other model.
+During MCT we observed that after one model reaches 100% training accuracy, its validation accuracy can still improve by learning to label for the other model.
 Possibly due to the rapid overfitting due to the increased view size (and consequently increased parameter count) the 1% split did not show improvement for either model.
+4.3 Experiments on Additional Datasets
+We compare MCT and CT to existing SoTA approaches that leverage unlabeled data directly during training (e.g., not just part of a pretext task).
+These experiments support our hypothesis that the primary cause of the failure of CT is that one or more of the views is sufficient to learn a model that achieves high accuracy.
+The closer to equivalent in performance individual views are, the less performance suffers.
+When the two views perform similarly and well, the performance improves.
+Tables 9 and 10 provide more information on these additional datasets.
+Furthermore, Figure 4 provides an example where CT performs as expected and accuracy increases across CT iterations.
+In all but one case MCT outperforms the SoTA top-1 accuracy.
+The main takeaway of these additional experiments, apart from establishing new SoTA results in datasets beyond ImageNet, is that MCT is in general more robust to view imbalance and provides better results than CT.
+As the performance degradation of co-training beyond the initial warmup period was not something expected, in Appendix D we provide charts similar to the one in Figure 4, but for the other cases of datasets or presence of labels in the training set, thus, giving full details on the results that we obtained for CT.
+Table 9: Additional comparisons to REACT.
+The authors include experiments for zero-shot performance and 10% of available labels.
+Results shown are for 10% of labels.
+For Flowers102 this is 1-shot performance.
 Dataset
@@
 Co-Training (MLP) 99.2 94.7 36.4
-MCT (MLP) 99.6 94.8 40.1
-Table 7: Additional Comparisons to REACT [52].
-The authors include experiments for zero-shot performance and 10% of available labels.
-Results above are for 10% of labels.
-For Flowers102 this is 1-shot performance.
+Meta Co-Training (MLP) 99.6 94.8 40.1
+Table 10: Additional comparisons to Semi-ViT using 10% (1%) of the available labels for training. For the iNaturalist task we use only the 1010 most frequent classes.
 Dataset
@@
 Co-Training (MLP) 59.7 (29.5) 94.7 (83.9)
-MCT (MLP) 76.0 (58.1) 94.8 (91.7)
-Table 8: Additional Comparisons to Semi-ViT using 10% (1%) of the available labels for training. For the iNaturalist task we use only the 1010 most frequent classes as in [17].
-Additional Datasets.
-In Tables 7 and 8 we compare meta co-training and co-training to existing SotA approaches that leverage unlabeled data directly during training (e.g., not just part of a pretext task).
-These experiments support our hypothesis that the primary cause of the failure of co-training is the insufficiency of one or more views.
+Meta Co-Training (MLP) 76.0 (58.1) 94.8 (91.7)
+Figure 4: The top-1 accuracy of the CT predictions for each iteration of CT.
+10% of available Food101 labels are used for training.
+The method exhibits performance improvement over multiple iterations of pseudo-labeling and retraining.
+5 Related Work
+While SSL has become very relevant in recent years with the influx of the data outpacing the availability of human labels, the paradigm and its usefulness have been studied at least since the 1960’s; e.g., [25, 16, 48, 22].
+Among the main approaches for SSL that are not directly related to our work, one popular direction encourages entropy minimization and consistency regularization using data augmentation such as MixMatch [4], UDA [51], ReMixMatch [3], FixMatch [49], and FlexMatch [53].
+Self-Training and Student-Teacher Frameworks.
+One of the earliest ideas for SSL is that of self-training [48, 22].
+This idea is still very popular with different student-teacher frameworks [55, 52, 43, 35, 24, 9].
+This technique is also frequently referred to as pseudo-labeling [33] as the student tries to improve its performance by making predictions on the unlabeled data, then integrating the most confident predictions to the labeled set, and subsequently retraining using the broader labeled set.
+In addition, one can combine ideas; e.g., self-training with clustering [30], or create ensembles via self-training [31], potentially using graph-based information [38].
+An issue with these approaches is that of confirmation bias [1], where the student cannot improve a lot due to incorrect labeling that occurs either because of self-training, or due to a fallible teacher.
+Co-training Algorithms.
+Co-training was introduced by Blum and Mitchell [5] and their paradigmatic example was classifying web pages using two different views: (i) the bags of words based on the content that the webpages had, and (ii) the bag of words formed by the words used in hyperlinks pointing to these webpages.
+In our context the two different views are obtained via different embeddings that we obtain when we use images as inputs to different foundation models and observe the resulting embeddings.
+Co-training has been very successful with wide applicability [20] and can outperform learning algorithms that use only single-view data [40].
+Furthermore, the generalization ability of co-training is well-motivated theoretically [5, 15, 2, 14].
+There are several natural extensions of the base idea, such as methods for constructing different views [50, 7, 11, 8], including connections to active learning [19], methods that are specific to deep learning [44], and other extensions that exploit more than two views on data; e.g., [54].
+6 Conclusion and Future Work
+We proposed MCT, a novel semi-supervised learning algorithm.
+We showed that our algorithm outperforms co-training in general and establishes a new SoTA top-1 classification accuracy on the ImageNet 10%.
+We also tested the performance of deep ensembles on ImageNet 10% and 1%.
+In both cases MCT achieved better results, demonstrating that MCT provides utility beyond what can be attributed to ensembling.
+In addition, using MCT, we were able to achieve new SoTA few-shot top-1 classification accuracy on other popular CV benchmark datasets, including one-shot top-1 classification accuracy for the Flowers102 dataset.
+Furthermore, we presented evidence that self-supervised learned representations are good candidates for views.
+We showed that the views constructed this way likely contain independent information and that models trained on these views, or concatenation of these views, can complement each other to produce stronger classifiers.
+Our experiments indicate that our method performs best when it has access to enough labeled data to produce strong views for pseudo-labeling.
+For some datasets this may preclude its application in one-shot settings.
+Further, if no sufficient and independent views are available and none can be constructed then we do not expect our model to perform meaningfully better than MPL.
+For future work we expect that fine-tuning the representation learners that we used for CT and MCT would improve results.
+Papers proposing foundation models observe performance benefits when fine-tuning those models on downstream tasks.
+Furthermore, we did not make use of the large available unlabeled datasets for CV in this study.
+Giving MCT access to a large repository of unlabeled data, as well as the entirety of the labeled data, should result in a performance boost.
+Finally, our list of pretext tasks was not exhaustive.
+It is an interesting question for future work which other tasks and models might yield useful views.
+7 Acknowledgments
+This material is based upon work supported by the U.S.
+National Science Foundation under Grant No.
+RISE-2019758.
+This work is part of the NSF AI Institute for Research on Trustworthy AI in Weather, Climate, and Coastal Oceanography (NSF AI2ES).
+Some of the computing for this project was performed at the OU Supercomputing Center for Education & Research (OSCER) at the University of Oklahoma (OU).
+Appendix A Further Details on Related Methods
+Our work is inspired by two SSL methods: meta pseudo labels (MPL) [43] and co-training [5].
+We provide more information on these methods below.
+This discussion should provide further insight for Algorithm 1.
+Meta Pseudo Labels.
+In order to discuss Meta Pseudo Labels, first we need to discuss Pseudo Labels.
+Pseudo Labels (PL) [33] is a broad paradigm for SSL, which involves a teacher and a student in the learning process.
+The teacher looking at the unlabeled set $U$, selects a few instances and assigns pseudo-labels to them.
+Subsequently these pseudo-examples are shown to the student as if they were legitimate examples so that the student can learn a better model.
+In this context, the teacher $f_{\theta_{T}}$ is fixed and the student is trying to learn a better model by aligning its predictions to pseudo-labeled batches $(X_{u},f_{\theta_{T}}(X_{u}))$.
+Thus, on a pseudo-labeled batch, the student optimizes its parameters using
+$\theta_{S}^{\text{PL}}\in\operatorname*{arg\,min}_{\theta_{S}}\ell\left(\operatorname*{arg\,max}_{\xi}f_{\theta_{T}}(X_{u})|_{\xi},f_{\theta_{S}}(X_{u})\right)\,.$ (7)
+The hope is that updated student parameters $\theta_{S}^{\text{PL}}$ will behave well on the labeled data $L=(X_{L},Y_{L})=\left\{(x_{i},y_{i})\right\}_{i=1}^{m}$; that is, the loss $\mathcal{L}_{L}\left(\theta_{S}^{\text{PL}}\left(\theta_{T}\right)\right)=\ell\left(Y_{L},f_{\theta_{S}^{\text{PL}}}(X_{L})\right)$ is small, where $\mathcal{L}\left(\theta_{S}^{\text{PL}}\left(\theta_{T}\right)\right)$ is defined so that the dependence between $\theta_{S}$ and $\theta_{T}$ is explicit.
+Meta Pseudo Labels (MPL) [43] is exploiting this dependence between $\theta_{S}$ and $\theta_{T}$.
+In other words, because the loss that the student suffers on $\mathcal{L}_{L}\left(\theta_{S}^{\text{PL}}\left(\theta_{T}\right)\right)$ is a function of $\theta_{T}$, and making a similar notational convention for the unlabeled batch $\mathcal{L}_{u}\left(\theta_{T},\theta_{S}\right)=\ell\left(\operatorname*{arg\,max}_{\xi}f_{\theta_{T}}(X_{u})|_{\xi},f_{\theta_{S}}(X_{u})\right)$, then one can further optimize $\theta_{T}$ as a function of the performance of $\theta_{S}$:
+$\left\{\begin{array}[]{rl}\displaystyle\min_{\theta_{T}}&\displaystyle\mathcal{L}_{L}\left(\theta_{S}^{\text{PL}}\left(\theta_{T}\right)\right)\,,\\[3.00003pt] \text{where}&\displaystyle\theta_{S}^{\text{PL}}(\theta_{T})\in\operatorname*{arg\,min}_{\theta_{S}}\mathcal{L}_{u}\left(\theta_{T},\theta_{S}\right)\,.\end{array}\right.$ (8)
+How we might exploit the dependency between $\theta_{T}$ and $\theta_{S}$ might not be immediately obvious.
+Eventually the models are optimized using $\theta_{S}^{\text{PL}}\approx\theta_{S}-\eta_{S}\cdot\nabla_{\theta_{S}}\mathcal{L}_{u}\left(\theta_{T},\theta_{S}\right)$ as a practical approximation [36, 21] which leads to the tractable teacher objective:
+$\min_{\theta_{T}}\mathcal{L}_{L}\left(\theta_{S}-\eta_{S}\cdot\nabla_{\theta_{S}}\mathcal{L}_{u}\left(\theta_{T},\theta_{S}\right)\right)\,.$ (9)
+For a more detailed explanation of this relationship we refer the interested reader to [43].
+Co-Training.
+Co-training relies on the existence of two different views of the instance space $\mathcal{X}$; i.e., $\mathcal{X}=\mathcal{X}^{(1)}\times\mathcal{X}^{(2)}$.
+Hence, an instance $x\in\mathcal{X}$ has the form $x=(x^{(1)},x^{(2)})\in\mathcal{X}^{(1)}\times\mathcal{X}^{(2)}$.
+We call the views $x^{(1)}$ and $x^{(2)}$ complementary.
+Co-training requires two assumptions in order to work.
+First, it is assumed that, provided enough many examples, each separate view is enough for learning a meaningful model.
+The second is the conditional independence assumption [5], which states that $x^{(1)}$ and $x^{(2)}$ are conditionally independent given the label of the instance $x=(x^{(1)},x^{(2)})$.
+Thus, with co-training, two models $f^{(1)}$ and $f^{(2)}$ are learnt, which help each other using different information captured from the different views of each instance.
+Below we provide more information.
+Given a dataset $S=L\cup U$ (Section 2 has notation), we can have access to two different views of the dataset due to the natural partitioning of the instance space $\mathcal{X}$; that is, $S^{(1)}=L^{(1)}\cup U^{(1)}$ and $S^{(2)}=L^{(2)}\cup U^{(2)}$.
+Initially two models $f^{(1)}$ and $f^{(2)}$ are learnt using a supervised learning algorithm based on the labeled sets $L^{(1)}$ and $L^{(2)}$ respectively.
+Learning proceeds in iterations, where in each iteration a number of unlabeled instances are integrated with the labeled set by assigning pseudo-labels to them that are predicted by $f^{(1)}$ and $f^{(2)}$, so that improved models $f^{(1)}$ and $f^{(2)}$ can be learnt by retraining on the augmented labeled sets.
+In particular, in each iteration, each model $f^{(v)}$ predicts a label for each view $x^{(v)}\in U^{(v)}$.
+The top-$k$ most confident predictions by model $f^{(v)}$ are used to provide a pseudo-label for the respective complementary instances and then these pseudo-labeled examples are integrated in to the labeled set for the complementary view.
+In the next iteration a new model is trained based on the augmented labeled set.
+Both views of the used instances are dropped from $U$ even if only one of the views may be used for augmenting the appropriate labeled set.
+This process repeats until the unlabeled sets $U^{(v)}$ are exhausted, yielding models $f^{(1)}$ and $f^{(2)}$ corresponding to the two views.
+Co-Training is evaluated using the joint prediction of these models which is the re-normalized element-wise product of the two predictions:
+$\frac{f^{(1)}\left(x^{(1)}\right)\otimes f^{(2)}\left(x^{(2)}\right)}{\left\lVert f^{(1)}\left(x^{(1)}\right)\otimes f^{(2)}\left(x^{(2)}\right)\right\rVert_{1}}\,.$ (10)
+where $\otimes$ is the Hadamard (element-wise) product of two vectors; that is, for vectors $\alpha,\beta\in\mathbb{R}^{d}$ their Hadamard (element-wise) product is the vector $\gamma=\alpha\otimes\beta$, where $\gamma_{j}=\alpha_{j}\cdot\beta_{j}$ for $j\in\{1,\ldots,d\}$.
+Appendix B Hyperparameter Configuration
+The relevant hyperparameters for training are included in Table 11 for co-training (CT), meta co-training (MCT), and deep ensembles.
+Table 11: Hyperparmeter configuration for CT, MCT, and deep ensemble. Other models used (MLP, linear) are trained for the same number of steps as MCT, with the same optimization parameters.
+config Co-Training Meta Co-Training Deep Ensemble
+optimizer Adam Adam Adam
+optimizer momentum $\beta_{1},\beta_{2}=0.9,0.999$ $\beta_{1},\beta_{2}=0.9,0.999$ $\beta_{1},\beta_{2}=0.9,0.999$
+base learning rate 1e-4 1e-4 1e-4
+batch size 4096 4096 4096
+learning rate schedule Reduce On Plateau Reduce On Plateau Reduce On Plateau
+schedule patience 10 steps 10 steps 10 steps
+decay factor 0.5 0.5 0.5
+training steps 200 (per iteration) 1000 1000 (per model)
+warmup steps n/a 200 n/a
+Appendix C Omitted Experimental Analysis on ImageNet
+In this section we provide some figures illustrating the experimental results on ImageNet 10% and ImageNet 1% that we did not include into the main text due to space limitations.
+In Figures 5 and 6 are shown results comparable to those in Figures 2 and 3 but for the ImageNet-1% dataset.
+As a reminder these figures are showing top-1 accuracy for CT and MCT when the constituent views used in the embeddings are CLIP and DINOv2.
+Similar to the case of ImageNet 10% we see that the CT performance decreases after the supervised warmup.
+Unlike CT, MCT performance increases after the supervised warmup.
+Figure 5: Top-1 accuracy of CT iterations on the CLIP and DINOv2 views for the ImageNet 1% dataset.
+Figure 6: MCT using the CLIP and DINOv2 views as a function of the training step. Models are trained on 1% of the ImageNet labels.
+Figure 7: Top-1 accuracy of CT iterations for the ImageNet 10% dataset. The average over all 10 pairs, the best pair, and the worst pair at each iteration are shown.
+Figure 8: Aggregate statistics of MCT across all views.
+Models are trained on 10% of the ImageNet labels.
+The maximum performance of any pair, the minimum performance of any pair, and the average of all 10 pairs is shown.
+Figure 9: Top-1 accuracy of CT iterations for the ImageNet 1% dataset. The average over all 10 pairs, the best pair, and the worst pair at each iteration are shown.
+Figure 10: Aggregate statistics of MCT across all views.
+Models are trained on 1% of the ImageNet labels.
+The maximum performance of any pair, the minimum performance of any pair, and the average of all 10 pairs is shown.
+In Figures 7 and 8 we show the min, max, and mean performance of all of the ten pairs of views for the ImageNet-10% dataset.
+Note that these are not three runs, but aggregate statistics over all runs, so the minimum and maximum points on the graph for one step may not correspond to the same pair as another step.
+Figures 9 and 10 are comparable to Figures 7 and 8, but for the ImageNet-1% dataset.
+Appendix D Further Details on Co-Training Iterations on Additional Datasets
+In this section we provide further details on the performance of pure CT on additional datasets, using the two views that appear to have the highest information content among those that we tried; i.e., CLIP and DINOv2.
+These experiments support our hypothesis that the primary cause of the failure of CT is that one or more views can be weak and do not suffice to learn a model that achieves reasonably high accuracy.
 The closer to equivalent in performance individual views are, the less the performance suffers.
 When the two views perform similarly and well, the performance improves.
-Section A in the appendix provides more information on these additional datasets; Figure 9 provides an example where co-training performs as expected.
-Meta co-training is in general more robust to view imbalance and provides better results.
-6 Conclusion and Future Work
-We presented a novel method of view construction for co-training approaches based on pretext tasks.
-We showed that the views themselves likely contain independent information and that models trained on these views can complement each other to produce stronger classifiers.
-However, these views appear to be insufficient for co-training on the ImageNet task.
-We proposed meta co-training, a novel algorithm for co-training to address the challenges faced by the co-training algorithm.
-We showed that our algorithm outperforms co-training in general, and establishes a new SotA top-1 classification accuracy on the ImageNet-10% dataset as well as new SotA few-shot top-1 classification accuracy on several other fine-grained image classification datasets.
-For future work we expect that fine-tuning the representation learners that we used for co-training and meta co-training would yield better results; which should be in line to observations from papers proposing foundation models that report performance benefits of fine-tuning those models on downstream tasks.
-Furthermore, we did not make use of the large available unlabeled datasets for computer vision in this study.
-We suspect that giving our models access to a large repository of unlabeled data as well as the entirety of the labeled data for any dataset on which we performed experiments would result in a performance boost.
-Finally, our list of pretext tasks was not exhaustive.
-It is an interesting question for future work which other tasks and models might yield useful views.
-Acknowledgements.
-This material is based upon work supported by the National Science Foundation under Grant No.
-ICER-2019758.
-This work is part of the NSF AI Institute for Research on Trustworthy AI in Weather, Climate, and Coastal Oceanography (AI2ES).
-Some of the computing for this project was performed at the OU Supercomputing Center for Education & Research (OSCER) at the University of Oklahoma (OU).
-Appendix A Further Analysis of Meta Co-Training and Co-Training
-Figures 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, and 15 include the top-1 accuracy of co-training using the best (CLIP and DINO) views found for the ImageNet task.
-In all cases except Figure 9, co-training reaches its maximum performance in the first iteration.
-Note that there is no 1% run for Flowers102 because there are only 10 examples for each class in the training set (see the discussion in Section B).
-A.1 Further Analysis of Meta Co-Training Experiments on ImageNet
-In Figures 5 and 6 are shown aggregate results comparable to those in Figures 1 and 2 in the main text, but for meta co-training.
-The min, max, and mean performance of any of the ten pairs is shown.
-Note that these are not three runs, but aggregate statistics over all runs, so the minimum and maximum points on the graph for one step may not correspond to the same pair as another step.
-Figure 5: Aggregate statistics of meta co-training across all views.
-Models are trained on 10% of the ImageNet labels.
-The maximum performance of any pair, the minimum performance of any pair, and the average of all 10 pairs is shown.
-Figure 6: Aggregate statistics of meta co-training across all views.
-Models are trained on 1% of the ImageNet labels.
-The maximum performance of any pair, the minimum performance of any pair, and the average of all 10 pairs is shown.
-In Figures 7 and 8 are shown results comparable to those in Figures 3 and  4 in the main text, but for meta co-training.
-Unlike co-training, meta co-training performance increases after the supervised warmup.
-Figure 7: Meta co-training using the CLIP and DINOv2 views as a function of the training step. Models are trained on 10% of the ImageNet labels.
-Figure 8: Meta co-training using the CLIP and DINOv2 views as a function of the training step. Models are trained on 1% of the ImageNet labels.
-A.2 Co-Training and Meta Co-Training on Food101
-In Figures 9 and  10 we show the top-1 accuracy of the iterations of co-training on the Food101 classification task.
-Co-training performs as expected with the pseudo labels allowing the other model to improve in the 10% case, but in the 1% case it appears that one or more of the views was insufficient and the performance decreases rapidly after the first iteration.
-Co-training achieves top-1 accuracy of 94.7% (resp., 83.9%) when 10% (resp., 1%) of the available labels for training are used from the Food101 dataset.
-Figure 9: The top-1 accuracy of the co-training predictions for each iteration of co-training.
-10% of available Food101 labels are used for training.
+Figure 11 and Figure 12 show CT iteration performance on the FGVC Aircraft dataset using 10% and 1% of labels respectively.
+Figure 11: The top-1 accuracy of the CT predictions for each iteration of CT.
+10% of available FGVC Aircraft labels are used for training.
 The method exhibits performance improvement over multiple iterations of pseudo-labeling and retraining.
-Figure 10: The top-1 accuracy of the co-training predictions for each iteration of co-training. 1% of available Food101 labels are used for training.
-Similarly, meta co-training achieves top-1 accuracy of 94.8% (resp., 91.7%) when 10% (resp., 1%) of the available training examples of the Food101 dataset are used.
-These results are compared against each other as well as against REACT and Semi-ViT in Tables 7 and 8 respectively.
-A.3 Co-Training and Meta Co-Training on iNaturalist
-In Figures 11 and 12 we show the top-1 accuracy of the iterations of co-training on the iNaturalist dataset.
-In both splits we see a steady decline in top-1 accuracy after the first iteration.
-Co-training achieves top-1 accuracy of 59.7% (resp., 29.5%) when only 10% (resp., 1%) of the training examples of the iNaturalist dataset are used.
-Figure 11: The top-1 accuracy of the co-training predictions for each iteration of co-training.
+Figure 12: The top-1 accuracy of the CT predictions for each iteration of CT. 1% of available FGVC Aircraft labels are used for training.
+Neither view performs well and there is a large difference between the performance of each view. The performance of CT quickly collapses.
+Figure 4, which was already presented in the main body of this work, provides an example where CT performs as expected and accuracy increases as CT progresses, though it can be seen from Figure 13 that the degeneration still occurs in the case with less labels.
+In Figure 13 we show the iterations of CT for Food101-1%.
+Figure 4 with the iterations for Food101-10% appears in 4.
+The 1-shot results for Flowers-102 (10%) appear in Figure 14.
+The iNaturalist-2017 results for 10% and 1% respectively appear in Figures 15, 16.
+Figure 13: The top-1 accuracy of the CT predictions for each iteration of CT. 1% of available Food101 labels are used for training.
+Figure 14: The top-1 accuracy of the CT predictions for each iteration of CT. 10% of available Flowers102 labels are used for training.
+Figure 15: The top-1 accuracy of the CT predictions for each iteration of CT.
 Only the 1010 most frequent classes are included.
 10% of available iNaturalist labels from those classes are used for training.
-Figure 12: The top-1 accuracy of the co-training predictions for each iteration of co-training.
+Figure 16: The top-1 accuracy of the CT predictions for each iteration of CT.
 Only the 1010 most frequent classes are included.
 1% of available iNaturalist labels from those classes are used for training.
-Along these lines meta co-training achieves 76.0% (resp., 58.1%) when only 10% (resp., 1%) of the training examples of the iNaturalist dataset are used.
-These results are compared against each other as well as against REACT and Semi-ViT in Tables 7 and 8 respectively.
-A.4 Co-Training and Meta Co-Training on CTFGVCAircraft
-In Figures 13 and 14 we show the top-1 accuracy of the iterations of co-training on the CTFGVCAircraft dataset.
-In both cases both views have low accuracy, so it is unsurprising that in subsequent iterations the accuracy decreases rapidly.
-In particular, co-training achieves top-1 accuracy of 36.4% (resp., 21.6%) when only 10% (resp., 1%) of the training examples of the CTFGVCAircraft dataset are used.
-Figure 13: The top-1 accuracy of the co-training predictions for each iteration of co-training. 10% of available FGVCAircraft labels are used for training.
-Figure 14: The top-1 accuracy of the co-training predictions for each iteration of co-training. 1% of available FGVCAircraft labels are used for training.
-Along these lines meta co-training achieves 40.1% (resp., 21.5%) when only 10% (resp., 1%) of the training examples of the CTFGVCAircraft dataset are used.
-The results corresponding to the 10% case of co-training and meta co-training are compared against REACT in Table 7, since the authors of [52] report results on zero-shot and 10% cases only.
-However, for completeness, we have decided to include Figure 14 which corresponds to the experiments that we did using only 1% of the labeled training data available for the CTFGVCAircraft dataset.
-Dataset #train #test #classes
-FGVCAircraft $3$ $3$ 100
-Flowers102 1020 6149 102
-Food101 75,750 25,250 101
-iNaturalist 175,489 29,083 1010
-ImageNet 1,281,167 50,000 1000
-Table 9: Sizes of datasets used.
-Subsets drawn for training are always drawn in a stratified way to maintain the class distribution of the original data.
-For ImageNet we use the split published with the SimCLRv2 [24] repository.
-config MCT Co-Training
-optimizer Adam Adam
-optimizer momentum $\beta_{1},\beta_{2}=0.9,0.999$ $\beta_{1},\beta_{2}=0.9,0.999$
-base learning rate 1e-4 1e-4
-batch size 4096 4096
-learning rate schedule Reduce On Plateau Reduce On Plateau
-schedule patience 10 steps 10 steps
-decay factor 0.5 0.5
-training steps 1000 200 (per iteration)
-warmup steps 200 n/a
-Table 10: Hyperparmeter configuration for meta co-training and co-training.
-Other models used (MLP, linear) are trained for the same number of steps as meta co-training, with the same optimization parameters.
-A.5 Co-Training and Meta Co-Training on Flowers102
-As explained earlier, the Flowers102 dataset has only 10 training examples for each class and hence 10% corresponds to 1-shot learning.
-In this context, in Figure 15 we give the performance of the iterations of co-training on the Flowers102 dataset.
-Even though the DINO view performs very well in the first iteration, there is a large gap between the performance of the CLIP view and the DINO view which leads to decreased performance in subsequent iterations.
-Co-training achieves top-1 accuracy of 99.2% when only 10% of the training examples of the Flowers102 dataset are used for training (i.e., a single labeled image per class).
-In the same context of 1-shot learning, meta co-training achieves top-1 accuracy of 99.6%.
-Both of the above results are compared against each other and against REACT in Table 7.
-Figure 15: The top-1 accuracy of the co-training predictions for each iteration of co-training. 10% of available Flowers102 labels are used for training.
-Appendix B Further Information on the Datasets and the Splits
-Table 9 includes a description of the different sizes of the datasets used for training and testing, as well as the number of cases for each dataset.
-Note that in the 10% experiments on the Flowers102 dataset each class has only 1 label.
-All datasets are approximately class-balanced with the exception of the iNaturalist dataset.
-In all cases we maintain the original class distribution when sampling subsets.
-All subsets are created with seeded randomness of a common seed (13) which ensures a fair comparison between methods.
-Appendix C Hyperparameter Configuration
-The relevant hyperparameters for training are included in Table 10 for co-training and meta co-training (MCT).
+Main Takeaway.
+The main takeaway of these additional experiments, apart from establishing new SoTA results in datasets beyond ImageNet, is that MCT is in general more robust to view imbalance and provides better results than CT.